### v1

# Zero forcing irredundant sets

Abstract
Irredundance has been studied in the context of dominating sets, via the concept of private neighbor. Here irredundance of zero forcing sets is introduced via the concept of a private fort and the upper and lower zero forcing irrdedundance numbers $\operatorname{ZIR}(G)$ and $\operatorname{zir}(G)$ are defined. Bounds on $\operatorname{ZIR}(G)$ and $\operatorname{zir}(G)$ are established and graphs having extreme values of $\operatorname{ZIR}(G)$ and $\operatorname{zir}(G)$ are characterized. The effect of the join and corona operations is studied. As the concept of a zero forcing irrdedundant set is new, there are many questions for future research.
Keywords zero forcing; irredundance; ZIr-set; ZIR
AMS subject classification 05C69, 05C50, 05C57
1 Introduction
A set $S$ is a dominating set of a graph $G$ if every vertex of $G$ is in $S$ or a neighbor of a vertex in $S$. A set $S$ of vertices in a graph G is irredundant if each vertex in $S$ dominates a vertex of G that is not dominated by any other vertex in $S$ (this may be a neighbor or itself), called a private neighbor. The study of irredundance was introduced in 1978 by Cockayne, Hedetniemi and Miller [11] in relation to the study of minimal dominating sets. The upper and lower irredundance numbers are the maximum cardinality and minimum cardinality of a maximal irredundant set over all maximal irredundant sets. Zero forcing was introduced as a bound for the maximum nullity among real symmetric matrices whose pattern of nonzero off-diagonal entries is described by the edges of a graph in [2] and independently in other applications. We introduce zero forcing irredundant sets with a private fort playing the role of a private neighbor; the upper and lower zero forcing irredundance numbers are the maximum cardinality and minimum cardinality of a maximal zero forcing irredundant set over all maximal zero forcing irredundant sets. A graph $G=(V(G),E(G))$ is simple, undirected, finite, and $V(G)\neq\emptyset$.
Zero forcing is a propagation process on a graph $G$. Starting with an initial set of blue vertices, the process colors vertices blue by repeated applications of the color change rule: A blue vertex $u$ can change the color of a white vertex $w$ to blue if $w$ is the only white neighbor of $u$. A zero forcing set for $G$ is a subset of vertices $B$ such that the final coloring of $B$ is $V(G)$ (the final coloring, which is unique, is the result of applying the color change rule until no more changes are possible). The zero forcing number of $G$ is defined to be the minimum cardinality of a zero forcing set. Recently minimal zero forcing sets that are not necessarily minimum have been studied in [7]. The upper zero forcing number $\overline{\operatorname{Z}}(G)$ of a graph $G$ is the maximum cardinality of a minimal zero forcing set of $G$.
Let $G$ be a graph. A nonempty set $F\subseteq V(G)$ is a fort if for all $v\in V(G)\setminus F$, $|F\cap N(v)|\neq 1$. Note that $V(G)$ is a fort (vacuously) and in a connected graph of order at least two, a fort must contain at least two vertices. Forts present obstructions to zero forcing: Given a set $B$ of blue vertices, some vertex in $B$ can perform a force in $G$ if and only if $V(G)\setminus B$ is not a fort.
If $F$ is a fort and $B_{0}\cap F=\emptyset$, then $B_{0}$ is not a zero forcing set (because $B_{0}\subseteq V(G)\setminus F$ and $V(G)\setminus F$ is not a zero forcing set). Conversely, if $B_{0}\cap F\neq\emptyset$ for all forts $F$, then $B_{0}$ can perform a force. Since the property that $B_{0}\cap F\neq\emptyset$ for all forts $F$ is inherited by all supersets of $B_{0}$, this process can be repeated until all vertices are blue, which shows that $B_{0}$ is a zero forcing set. This equivalence, stated in the next theorem, was first noted by Brimkov, Fast and Hicks in [8].
Theorem 1.1.
[8] Let $G$ be a graph. Then $B\subseteq V(G)$ is a zero forcing set if and only if $B$ intersects every fort.
Vertices $v$ and $u$ of $G$ are neighbors if $vu$ is an edge of $G$, the (open) neighborhood of $v$ is $N_{G}(v)=\{u\in V(G):u\mbox{ and }v\mbox{ are neighbors}\}$, the closed neighborhood of $v$ is $N_{G}[v]=N_{G}(v)\cup\{v\}$, and $N_{G}[S]=\cup_{x\in S}N_{G}[x]$ for $S\subseteq V(G)$. When the graph $G$ is clear, open and closed neighborhoods can be denoted by $N(v),N[v],N[S]$.
Remark 1.2.
In general, a superset of a fort $F$ need not be a fort, because if vertex $v$ is added to $F$ there may be a vertex $u$ such that $u$ has no neighbors in $F$ but is adjacent to $v$. However, there are conditions under which all or certain supersets of a fort are a fort. If $F$ is a fort and $N(v)\subseteq F$, then $F\cup\{v\}$ is a fort. If $F$ is a fort such that $N(v)\subseteq F$ for every $v\not\in F$, then every superset of a fort is a fort. A union of forts is a fort: Suppose $F_{1}$ and $F_{2}$ are forts of $G$. If $u\not\in F_{1}\cup F_{2}$, then $u$ has zero or at least two neighbors in each of $F_{i}$, so $u$ has zero or at least two neighbors in $F_{1}\cup F_{2}$
A private fort is a natural parallel of a private neighbor.
Definition 1.3.
Let $G$ be a graph and let $S\subseteq V(G)$. For $x\in S$, a fort $F$ of $G$ is a private fort of $x$ (relative to $S$), if $S\cap F=\{x\}$. A minimal private fort of $x$ is a private fort of $x$ that does not properly contain another private fort of $x$.
Definition 1.4.
Let $G$ be a graph and let $S\subseteq V(G)$. Then $S$ is a $\operatorname{Z}$-irredundant set or ZIr-set if every element of $S$ has a private fort. The lower ZIr number is
[…]
The upper ZIr number is
[…]
We use the terms upper ZIR set and lower zir set to refer to maximal ZIr-sets $S$ and $S^{\prime}$ such that $|S|=\operatorname{ZIR}(G)$ and $|S^{\prime}|=\operatorname{zir}(G)$, respectively.
The domination number $\gamma(G)$ and the upper domination number $\Gamma(G)$ of $G$ are the minimum and maximum cardinalities of a minimal dominating set of $G$. Recall that $S\subseteq V(G)$ is an irredundant set if every vertex in $S$ has a private neighbor relative to $S$. The upper irredundance number $\operatorname{IR}(G)$ and the lower irredundance number $\operatorname{ir}(G)$ of $G$ are the maximum and minimum cardinalities of a maximal irredundant set of $G$.
A set that is both dominating and irredundant is a minimal dominating and maximal irredundant set. In fact, minimal dominating sets have been characterized as exactly those sets that are both dominating and irredundant [11]. We therefore have that $\operatorname{ir}(G)\leq\gamma(G)\leq\Gamma(G)\leq\operatorname{IR}(G)$.
[…]
Basic results about the relationship between $\operatorname{Z},\overline{\operatorname{Z}}$, $\operatorname{ZIR}$, and $\operatorname{zir}$ are presented in Section 2, with many parallels to $\gamma,\Gamma$, IR, and ir. Examples of upper ZIR and lower zir numbers for specific graph families are given in Section 3. Section 4 presents bounds on upper ZIR and lower zir numbers and graphs having extreme upper ZIR and lower zir numbers are characterized in Section 5. The effect of the join and corona operations is studied in Section 6 and it is shown in Section 7 that the lower zir number is noncomparable to several lower bounds for the zero forcing number. Finally, Section 8 presents a summary table of upper ZIR and lower zir numbers, zero forcing numbers and other related parameters for various graph families and discusses directions for future research.
[…]
Vertices $u$ and $w$ of a graph $G$ are twins if $N_{G}(u)=N_{G}(w)$ (independent twins) or $N_{G}[u]=N_{G}[w]$ (adjacent twins). A set of vertices $\{w_{1},\dots,w_{r}\}$ is a set of twins if every two vertices in the set are independent twins or every two vertices in the set are adjacent twins. The next result is well-known and useful in the study of zero forcing.
Proposition 1.5.
[15, Proposition 9.15] If a graph has a set of twins $\{w_{1},\dots,w_{r}\}$, then each zero forcing set must contain at least $r-1$ of these twins.
2 Preliminary results
Many of the results in this section are not difficult to prove but provide intuition and establish a collection of tools we shall frequently use. Since zero forcing is one of the primary inspirations for this work, we begin by investigating its relationship with ZIr-sets. We have an immediate parallel with dominating and irredundant sets.
Remark 2.1.
Let $S\subseteq V(G)$ be both a zero forcing set and ZIr-set. Then we can see that $S$ is a minimal zero forcing set and maximal ZIr-set: Since every element $x\in S$ has a private fort $F_{x}$, $(S\setminus\{x\})\cap F_{x}=\emptyset$ and $S\setminus\{x\}$ is not a zero forcing set. Since a superset of a zero forcing set is a zero forcing set, and no proper superset of $S$ is miminal, $S$ is a maximal ZIr-set.
Proposition 2.2.
Let $G$ be a graph and $S\subseteq V(G)$. Then $S$ is a minimal zero forcing set if and only if $S$ is a maximal ZIr-set of $G$ and $S\cap F\neq\emptyset$ for every fort $F$ of $G$.
Proof.
Suppose that $S$ is a minimal zero forcing set. Let $x\in S$. Since $S$ is minimal, $S\setminus\{x\}$ is not a zero forcing set. By Theorem 1.1 there exists a fort $F_{x}$ such that $(S\setminus\{x\})\cap F_{x}=\emptyset$ and $S\cap F_{x}\neq\emptyset$. Thus $F_{x}$ is a private fort of $x$ relative to $S$. This argument holds for every vertex in $S$ and so $S$ is a ZIr-set. Theorem 1.1 guarantees $S\cap F\neq\emptyset$ for every fort $F$ of $G$ and hence $S$ is a maximal ZIr-set.
Now suppose that $S$ is a maximal ZIr-set of $G$ and $S\cap F\neq\emptyset$ for every fort $F$ of $G$. Then Theorem 1.1 implies that $S$ is a zero forcing set, so $S$ is a minimal zero forcing set by Remark 2.1. ∎
Note that since $V(G)$ is a fort for any graph $G$, $1\leq\operatorname{zir}(G)$.
Corollary 2.3.
Let $G$ be a graph. Then $1\leq\operatorname{zir}(G)\leq\operatorname{Z}(G)\leq\overline{\operatorname{Z}}(G)\leq\operatorname{ZIR}(G)$.
Not every maximal ZIr-set is a zero forcing set, as the next example shows.
Example 2.4.
Consider the cycle $C_{5}$. The minimal forts are sets of three vertices that do not contain three consecutive vertices (consecutive in the order around the cycle). It follows that every set of two vertices is a maximal ZIr-set, but not every set of two vertices is a zero forcing set since it is well-known that a zero forcing set of $C_{5}$ must contain two adjacent vertices.
Next we present elementary results about ZIr sets. As the next remark illustrates, it often suffices to restrict to connected graphs when studying ZIr-sets.
Remark 2.5.
If $G=G_{1}\mathbin{\,\sqcup\,}\dots\mathbin{\,\sqcup\,}G_{k}$, then $S$ is a ZIr-set of $G$ if and only if $S\cap G_{i}$ is a ZIr-set of $G_{i}$ for $i=1,\dots,k$. Thus $\operatorname{zir}(G_{1}\mathbin{\,\sqcup\,}\dots\mathbin{\,\sqcup\,}G_{k})=\operatorname{zir}(G_{1})+\dots+\operatorname{zir}(G_{k})$ and $\operatorname{ZIR}(G_{1}\mathbin{\,\sqcup\,}\dots\mathbin{\,\sqcup\,}G_{k})=\operatorname{ZIR}(G_{1})+\dots+\operatorname{ZIR}(G_{k})$
The next example is an immediate consequence.
Example 2.6.
Let $n\geq 1$. Since $\operatorname{zir}(K_{1})=1$, $\operatorname{zir}(\overline{K_{n}})=Z(\overline{K_{n}})=\overline{Z}(\overline{K_{n}})=\operatorname{ZIR}(\overline{K_{n}})=n$.
Remark 2.7.
Suppose $G$ is a graph of order $n\geq 2$ that has an edge. Every fort in a connected component containing an edge must contain at least two vertices. Adding an additional private fort to a union of private forts must add a new vertex, so the union of all $k$ private forts associated with a maximal ZIr-set of $k$ vertices must contain at least $k+1$ vertices. Thus $\operatorname{ZIR}(G)\leq n-1$.
Remark 2.8.
Let $G$ be a graph with no isolated vertices. Then every minimal zero forcing set is a maximal ZIr-set and no vertex is in every minimal zero forcing set [3]. Thus there does not exist a vertex $v$ that is contained in every maximal ZIr-set of $G$.
Lemma 2.9.
Let $G$ be a graph with no isolated vertices and let $S\subseteq V(G)$. If $N[v]\subseteq S$ for some vertex $v$, then $S$ is not a ZIr-set, or equivalently, if $S$ is a ZIr-set of $G$, then $V(G)\setminus S$ is a dominating set of $G$.
Proof.
Suppose there exists a vertex $v\in S$ such that $N(v)\subseteq S$. Having assumed that $G$ has no isolated vertices, there exists a vertex $u\in N(v)$. Let $F\subseteq V(G)$ such that $F\cap S=\{u\}$. Since $N(v)\subseteq S$, $v$ is adjacent to exactly one element of $F$. Thus there does not exist a private fort of $u$ relative to $S$, and so $S$ is not a ZIr-set of $G$. ∎
It is well known (and easy to see) that $\delta(G)\leq\operatorname{Z}(G)$. The next two results strengthen this bound.
Proposition 2.10.
Let $G$ be a graph and $d\geq 1$. If $v_{1},\ldots,v_{d}\in V(G)$ all have degree at least $d$, then $S=\{v_{1},\ldots,v_{d}\}$ is a ZIr set.
Proof.
Assume $v_{1},\ldots,v_{d}\in V(G)$ all have degree at least $d$. It suffices to show that the set $F_{i}=(V(G)\setminus S)\cup\{v_{i}\}$ is a private fort of $v_{i}$ relative to $S$ for every $i=1,\ldots,d$. This is certainly the case if $d=1$ since $V(G)$ is a fort, so suppose $d\geq 2$. Let $v\in V(G)\setminus F_{i}$ for some $i=1,\ldots,d$. If $v$ has two or more neighbors in $V(G)\setminus S$, then $|N(v)\cap F_{i}|\geq 2$. So, assume $v$ has at most one neighbor in $V(G)\setminus S$. Since $v\in S$ and $|N(v)|\geq d$, $v$ is adjacent to exactly one vertex in $V(G)\setminus S$ and to $v_{i}$ because $S\setminus\{v\}\subseteq N(v)$. Thus $|F_{i}\cap N(v)|\geq 2$ as required. ∎
Corollary 2.11.
If $G$ is a graph and $\delta(G)\geq 1$, then every set of size $\delta(G)$ is a ZIr-set. For every graph $G$, $\delta(G)\leq\operatorname{zir}(G)$.
The next example shows the bound in Corollary 2.11 is sharp.
Example 2.12.
Let $n\geq{2}$. Then $\operatorname{zir}(K_{n})=n-1=\operatorname{ZIR}(K_{n})$ by Corollary 2.11 (since $\delta(K_{n})=n-1$) and Remark 2.7.
Example 2.12 implies that $\operatorname{Z}(K_{n})=\overline{Z}(K_{n})=n-1$, but this is already known [2, 6]. In the next section we establish values of the parameters $\operatorname{zir},\operatorname{Z},\overline{\operatorname{Z}},$ and $\operatorname{ZIR}$ for additional graph families. These results show that both $\operatorname{zir}(G)=\operatorname{Z}(G)=\overline{\operatorname{Z}}(G)=\operatorname{ZIR}(G)$ and $\operatorname{zir}(G)<\operatorname{Z}(G)<\overline{\operatorname{Z}}(G)<\operatorname{ZIR}(G)$ are possible (Example 2.12 and Proposition 3.6).
Observation 2.13.
Let $S$ be a ZIr-set of a graph $G$, of order $n$, let $x_{i}\in S$ for $i=1,\dots,k$, and let $F_{i}$ be a private fort of $x_{i}$. Then $|S|\leq n-|\cup_{i=1}^{k}F_{i}|+k$.
3 Determining ZIR and zir for graph families
In this section we establish the values of the upper and lower ZIr numbers for various families of graphs. In doing so, we highlight various techniques that will be used throughout the paper. We also include known zero forcing numbers and upper zero forcing numbers in each of the following results as the connections between the parameters and the underlying graph is rather interesting. See Table 8.1 for a summary of the values of $\operatorname{zir},\operatorname{Z},\overline{\operatorname{Z}}$, and $\operatorname{ZIR}$ determined for various graphs in this and other sections.
Let $G$ be a graph on $n$ vertices. The next example illustrates that the gap between $\operatorname{zir}(G)$ and $\operatorname{Z}(G)$ (and hence also $\operatorname{ZIR}(G)$) can be as large as $n-3$. It is known that $\operatorname{Z}(K_{q,p})=q+p-2$ and $\overline{\operatorname{Z}}(K_{q,p})=q+p-2$ [2, 6].
Example 3.1.
Suppose $1\leq q\leq p$ and let $U=\{u_{1},\ldots,u_{q}\}$ and $W=\{w_{1},\dots,w_{p}\}$ be the partite sets of $K_{q,p}$. Any set $S$ that omits at least one vertex $u_{k}$ from $U$ and at least one vertex $w_{\ell}$ from $W$ is a ZIr-set, since then $\{u_{i},u_{k}\}$ is a private fort of $u_{i}$ and $\{w_{j},w_{\ell}\}$ is a private fort of $w_{j}$ (if $q=1$, the forts $\{u_{i},u_{k}\}$ do not exist). If $S$ is a ZIr-set containing $U$, then $S=U$ because a private fort of $u_{i}$ necessarily contains $V$ and similarly for $W\subseteq S$ (note that when $q=1$, the private fort of $u_{1}$ is $V(K_{1,p})$). Thus $\operatorname{Z}(K_{q,p})=\overline{\operatorname{Z}}(K_{q,p})=\operatorname{ZIR}(K_{q,p})=q+p-2$ and $\operatorname{zir}(K_{q,p})=q$.
Proposition 3.2.
Let $n\geq 4$. Then $\operatorname{zir}(C_{n})=\operatorname{Z}(C_{n})=\overline{\operatorname{Z}}(C_{n})=2$ and $\operatorname{ZIR}(C_{n})=\left\lfloor\frac{n}{2}\right\rfloor$.
Proof.
It is known that $Z(C_{n})=\overline{Z}(C_{n})=2$ [2, 6]. Together with Corollary 2.11, this implies $\operatorname{zir}(C_{n})=2$. Let $v_{1},\ldots,v_{n}$ denote the vertices of $C_{n}$, where $v_{n}v_{1}\in E(C_{n})$ and $v_{i}v_{i+1}\in E(C_{n})$ for $i=1,\ldots,n-1$. Let $S=\{v_{2i-1}:i=1,\ldots,\lfloor\frac{n}{2}\rfloor\}$ and $F_{i}=(V(C_{n})\setminus S)\cup\{v_{i}\}$. By construction, $F_{2i-1}$ is a private fort of $v_{2i-1}$ for $i=1,\dots,\lfloor\frac{n}{2}\rfloor$. Thus $S$ is a ZIr-set. Since $|S|=\lfloor\frac{n}{2}\rfloor$, $\operatorname{ZIR}(C_{n})\geq\left\lfloor\frac{n}{2}\right\rfloor$.
Since every subset of $V(C_{n})$ that does not contain a pair of consecutive vertices has cardinality at most $\lfloor\frac{n}{2}\rfloor$, it suffices to show that if $S$ is a ZIr-set of $C_{n}$ that contains a pair of adjacent vertices, then $|S|=2$.
Assume that $S$ is a ZIr-set of $C_{n}$ that contains a pair of adjacent vertices. Without loss of generality let $\{v_{1},v_{2}\}\subseteq S$. Let $F$ be a fort of $C_{n}$. Suppose, to obtain a contradiction, that $F\cap\{v_{1},v_{2}\}=\emptyset$. Then $v_{3}\notin F$ since otherwise $|N(v_{2})\cap F|=1$. Recursively applying this argument implies $v_{k}\notin F$ for $k=4,\ldots,n$ and hence $F=\emptyset$, a contradiction. Thus every fort of $C_{n}$ intersects $\{v_{1},v_{2}\}$ non-trivially. Since every element of $S$ must have a private fort, it follows that $|S|=2$. ∎
Observe that $C_{3}=K_{3}$, so the parameters are determined by Example 2.12 (and $\operatorname{ZIR}(K_{3})=2>\lfloor\frac{3}{2}\rfloor$).
Proposition 3.3.
For $n\geq 1$, $\operatorname{zir}(P_{n})=\operatorname{Z}(P_{n})=1$. For $n\geq 4$, $\overline{\operatorname{Z}}(P_{n})=2$. For $n\geq 5$, $\operatorname{ZIR}(P_{n})=\left\lfloor\frac{n-1}{2}\right\rfloor$.
Proof.
It is well-known that $Z(P_{n})=1$ for $n\geq 1$, which implies $\operatorname{zir}(P_{n})=1$. For $n\geq 4$, $\overline{Z}(P_{n})=2$ [6]. Assume $n\geq 5$. Let $v_{1},\dots,v_{n}$ denote the vertices of $P_{n}$, where $v_{i}v_{i+1}\in E(P_{n})$ for $i=1,\dots,n-1$. Just as in the proof of Proposition 3.2, if $S$ is a ZIr-set of $P_{n}$ that contains a pair of adjacent vertices, then $|S|\leq 2$. Observe that every fort in $P_{n}$ contains $v_{1}$ and $v_{n}$.
Let $S=\{v_{2i}:i=1,\ldots,\lfloor\frac{n-1}{2}\rfloor\}$ and $F_{i}=(V(P_{n})\setminus S)\cup\{v_{i}\}$. Note that $|S|=\lfloor\frac{n-1}{2}\rfloor$. By construction, $F_{2i}$ is a private fort of $v_{2i}$ for $i=1,\dots,\lfloor\frac{n-1}{2}\rfloor$. Thus $S$ is a ZIr-set. Moreover, $S$ is maximal since $|S|\geq 2$ and the addition of any vertex to $S$ would introduce a pair of consecutive vertices or a vertex of degree 1. Since any set that does not contain a pair of consecutive vertices and does not contain $v_{1}$ or $v_{n}$ has cardinality at most $\lfloor\frac{n-1}{2}\rfloor$ we conclude that $\operatorname{ZIR}(P_{n})=\lfloor\frac{n-1}{2}\rfloor$. ∎
Observation 3.4.
Observe that $P_{2}=K_{2}$ and $P_{3}=K_{1,2}$, so $\operatorname{ZIR}(P_{2})=1$ and $\operatorname{ZIR}(P_{3})=1$ from Examples 2.12 and 3.1. It can be shown computationally [13] that $\operatorname{ZIR}(P_{4})=2$.
For $k\geq 2$, the $k$th-friendship graph $\operatorname{Fr}(k)$ is the graph with $12k+$ 1 vertices and $3k$ edges constructed as the join of one vertex to k disjoint copies of $K_{2}$; $\operatorname{Fr}(3)$ is shown in Figure 3.1. It is known that $\operatorname{Z}(\operatorname{Fr}(k))=k+1$ [15, Theorem 9.5].
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{6}$$v_{5}$
Figure 3.1: The friendship graph $\operatorname{Fr}(3)$
Proposition 3.5.
For $k\geq 2$, $\operatorname{zir}(\operatorname{Fr}(k))=\operatorname{Z}(\operatorname{Fr}(k))=\overline{\operatorname{Z}}(\operatorname{Fr}(k))=\operatorname{ZIR}(\operatorname{Fr}(k))=k+1$.
Proof.
Let $k\geq 2$ and let $V(\operatorname{Fr}(k))=\{v_{0},\ldots,v_{2k}\}$, where $v_{0}$ is the unique vertex of degree $2k$ and $v_{j}$ is adjacent to $v_{j+1}$ for $j\equiv 1\mod 2$ and $1\leq j<2k$. We begin by examining the forts of $\operatorname{Fr}(k)$. For $i=1,\dots,k$, let $F_{i}=\{v_{2i-1},v_{2i}\}$ and observe that $F_{i}$ is a (minimal) fort. A union of forts of the form $F_{i}$ is a fort by Remark 1.2. There are $2^{k}$ (minimal) forts that contain $v_{0}$ and exactly one element of each $F_{i}$ for $i=1,\dots,k$; a superset of such a fort is a fort by Remark 1.2, and is called a center-fort. These are the only forts of $\operatorname{Fr}(k)$ because a fort that does not contain $v_{0}$ is a union of forts of the form $F_{i}$ and a fort that does contain $v_{0}$ is a center-fort.
There are only $k$ forts of the form $F_{i}$ so if $S$ is a ZIr-set and $|S|\geq k+1$, then one of the private forts relative to $S$ must be a center-fort $F^{\prime}$, which contains at least $k+1$ vertices. Thus $|V(\operatorname{Fr}(k))\setminus F^{\prime}|\leq k$ and $|S|\leq k+1$. This implies $\operatorname{ZIR}(\operatorname{Fr}(k))\leq k+1$.
Let $S$ be a ZIr-set. We show that $S$ is contained in a ZIr-set with cardinality $k+1$, which implies every maximal ZIr-set has cardinality $k+1$. Suppose that $S\cap F_{\ell}=\emptyset$ for some $\ell\in\{1,\dots,k\}$ and choose a set of private forts relative to $S$ for the vertices of $S$. Without loss of generality, we may assume that every chosen private fort that is not a center-fort is of the form $F_{i}$ for some $i\in\{1,\dots,k\}$. We show that $S^{\prime}=S\cup\{v_{2\ell-1}\}$ is also a ZIr-set with private fort $F_{\ell}$ for $v_{2\ell-1}$ and appropriate modifications of private forts of the vertices in $S$. Suppose that the chosen private fort $F$ for some $v_{j}\in S$ contains $v_{2\ell-1}$. Note that $F$ is a center-fort. Since $v_{j}\not\in F_{\ell}$, $F^{\prime}=\left(F\cup\{v_{2\ell}\}\right)\setminus\{v_{2\ell-1}\}$ is a private fort relative to $S^{\prime}$ for $v_{j}$. Thus $S$ is not maximal, and there is a ZIr-set $\widehat{S}$ such that $S\subseteq\widehat{S}$ and $\widehat{S}\cap F_{i}\neq\emptyset$ for $i=1,\dots,k$. If $F_{i}\subseteq\widehat{S}$ for some $i$, then $|\widehat{S}|\geq k+1$. So assume $|\widehat{S}\cap F_{i}|=1$ for $i=1,\dots,k$. If $v_{0}\in\widehat{S}$, then $|\widehat{S}|\geq k+1$. Suppose $v_{0}\not\in\widehat{S}$. For $\widehat{S}\cup\{v_{0}\}$, we have private forts $F_{i}$ for elements of $\widehat{S}$ and the center-fort that does not contain any element of $\widehat{S}$ for $v_{0}$. Thus $\widehat{S}\cup\{v_{0}\}$ is a ZIr-set and every maximal ZIr-set has cardinality at least $k+1$. ∎
For $r\geq 3$, denote the two vertices in the smaller partite set of $K_{2,r}$ by $u$ and $u^{\prime}$ and the set of vertices in the larger partite set by $W=\{w_{1},\dots,w_{r}\}$. For odd $s\geq 5$, denote the set of vertices of $P_{s}$ by $Y=\{y_{1},\dots y_{s}\}$ in path order. Let $H(r,s)$ be the graph of order $r+s+1$ constructed from $K_{2,r}$ and $P_{s}$ by identifying $u^{\prime}$ with $y_{s}$, retaining the label $y_{s}$ (see Figure 3.2). The next example shows that all the parameters $\operatorname{zir},\operatorname{Z},\overline{\operatorname{Z}},$ and $\operatorname{ZIR}$ can be different (for example, $\operatorname{zir}(H(3,5))=2$, $\operatorname{Z}(H(3,5))=3$, $\overline{\operatorname{Z}}(H(3,5))=4$ and $\operatorname{ZIR}(H(3,5))=5$ [13]).
$y_{1}$$y_{2}$$y_{3}$$y_{4}$$y_{5}$$w_{2}$$u$$w_{1}$$w_{3}$
Figure 3.2: The graph $H(3,5)$, which has different values for $\operatorname{zir},\operatorname{Z},\overline{\operatorname{Z}},$ and $\operatorname{ZIR}$
Proposition 3.6.
For $r\geq 2$ and odd $s\geq 3$, $\operatorname{zir}(H(r,s))=2$, $\operatorname{Z}(H(r,s))=r$, and $\operatorname{ZIR}(H(r,s))=r+\frac{s-1}{2}$. For $r\geq 2$ and odd $s\geq 5$, $\overline{\operatorname{Z}}(H(r,s))=r+1$. For $r\geq 2$, $\overline{\operatorname{Z}}(H(r,3))=r$.
Proof.
Let $r\geq 2$ and let $s$ be odd with $s\geq 3$. By Proposition 1.5, any zero forcing set must contain at least $r-1$ of the vertices $\{w_{1},\dots,w_{r}\}$ since this is a set of twins. The only minimal zero forcing sets are $r-1$ of these vertices together with exactly one of the following sets: $\{u\}$, $\{y_{1}\}$, $\{y_{s}\}$, or (if $s\geq 5$) $\{y_{i},y_{i+1}\}$ for $i=2,\dots,s-2$. Thus $\operatorname{Z}(H(r,s))=r$; $\overline{\operatorname{Z}}(H(r,s))=r+1$ if $s\geq 5$ and $\overline{\operatorname{Z}}(H(r,3))=r$.
If $F$ is a fort such that $F\cap Y\neq\emptyset$, then $y_{1}\in F$ and at least every other vertex of $Y$ is in $F$. For the ZIr-set $S_{1}=\{u,y_{1}\}$, the only private fort $F_{u}$ of $u$ is $\{u,w_{1},\dots,w_{r}\}$, so $S_{1}$ is maximal and $\operatorname{zir}(H(r,s))\leq 2$. Since $S_{1}$ is a ZIr-set, $\{y_{1}\}$ is not maximal. Since every other vertex has degree at least two, every other vertex can also be in a ZIr-set of two vertices by Proposition 2.10. Thus $\operatorname{zir}(H(r,s))=2$.
Let $D=\{y_{1},y_{3},\dots,y_{s-2},y_{s},u\}$ and $S_{2}=V(H(r,s))\setminus D$. Every vertex $v\in S_{2}$ is adjacent to at least two vertices in $D$, so $\{v\}\cup D$ is a private fort of $v$ and $S_{2}$ is a ZIr-set. This shows that $\operatorname{ZIR}(H(r,s)\geq r+\frac{s-1}{2}$.
Now let $S$ be a ZIr-set. There are $r+1$ vertices not in $Y$, and they cannot all be in $S$ because if $w_{j}\in S$ then a private fort of $w_{j}$ must contain $u$ or some $w_{k}$ with $k\neq j$. So $|S\cap\{u,w_{1},\dots,w_{r}\}|\leq r$. If $S$ does not contain two consecutive vertices in $Y$, then $|S\cap Y|\leq\frac{s-1}{2}$. So assume that $S$ contains two consecutive vertices $y_{i-1},y_{i}$. Then $S\cap Y\subseteq\{y_{i-1},y_{i}\}$ because a fort $F$ such that $F\cap Y\neq\emptyset$ cannot omit two consecutive vertices of $Y$. Thus $|S\cap Y|\leq 2\leq\frac{s-1}{2}$. In all cases, $|S|\leq r+\frac{s-1}{2}$
The value of $\operatorname{ZIR}$ is established for additional families of graphs in Theorem 4.4, 4.8 and 5.11 and Proposition 6.4 below.
4 Bounds
In this section we use known results about 2-domination to establish bounds on $\operatorname{ZIR}(G)$ and other methods to establish additional bounds on $\operatorname{ZIR}(G)$. A $2$-dominating set of a graph $G$ is a subset $D\subseteq V(G)$ such that every vertex not in $D$ is adjacent to at least two vertices in $D$. The minimum cardinality of a 2-dominating set is the $2$-domination number of the graph, indicated by $\gamma_{2}(G)$. The idea used in the next proof to show that $n-\gamma_{2}(G)\leq\operatorname{ZIR}(G)$ was already used informally in Proposition 3.6 and other results.
Proposition 4.1.
[…]
Proof.
Let $S$ be $\operatorname{ZIR}$-set of $G$ such that $|S|=\operatorname{ZIR}(G)$. By Lemma 2.9, $V(G)\setminus S$ is a dominating set and therefore $\gamma(G)\leq|V(G)|-|S|=n-\operatorname{ZIR}(G)$.
Let $D$ be a $\gamma_{2}$-set of $G$ and let $S=V(G)\setminus D=\{v_{1},\dots,v_{k}\}$. The set $F_{i}=D\cup\{v_{i}\}$ is a private fort of $v_{i}$ relative to $S$. Therefore $S$ is a ZIr-set and $\operatorname{ZIR}(G)\geq|S|$.
Example 3.1 shows $K_{q,p}$ with $q\geq 2$ realizes equality for the upper bound and Proposition 3.2 shows $C_{n}$ realizes equality for the lower bound. ∎
Note that the previous lower bound does not need the restriction that $G$ has no isolated vertices.
We use known results of Caro and Roddity [9] and Cockayne et al. [10] bounding $\gamma_{2}(G)$ for graphs with specified minimum degree to bound $\operatorname{ZIR}(G)$ for such graphs; note that the results in [9] and [10] are more general and imply Theorems 4.2 and 4.6.
Theorem 4.2.
[9] Let $G$ be a graph of order $n$. If $\delta(G)\geq 3$, then $\gamma_{2}(G)\leq\frac{n}{2}$.
Corollary 4.3.
Let $G$ be a graph of order $n\geq 2$ with $\delta(G)\geq 3$. Then $\operatorname{ZIR}(G)\geq\frac{n}{2}$.
This bound is sharp, as is shown in the next result. A diamond is a $K_{4}$ with one edge deleted. Let $D_{1},D_{2},\dots,D_{n}$ be $n$ diamonds each with vertex set $V_{i}=\{a_{i},b_{i},c_{i},d_{i}\}$ where the edge $a_{i}c_{i}$ is missing. Define the graph consisting of the $k$ diamonds together with the edges $c_{i}a_{i+1}$ for $11\leq i\leq k-$ 1 and $c_{k}a_{1}$ to be the $k$th necklace, denoted by $N_{k}$. The 3rd necklace $N_{3}$ is illustrated in Figure 4.1.
$a_{1}$$b_{1}$$c_{1}$$d_{1}$$a_{2}$$b_{2}$$c_{2}$$d_{2}$$a_{3}$$b_{3}$$c_{3}$$d_{3}$
Figure 4.1: The graph $N_{3}$
Theorem 4.4.
Let $k\geq 2$. Then $\operatorname{ZIR}(N_{k})=2k=\frac{1}{2}|V(N_{k})|$.
Proof.
By Corollary 4.3, $\operatorname{ZIR}(N_{k})\geq 2k$. Let $S$ be a ZIr-set of $N_{k}$. If $|V_{i}\cap S|\leq 2$ for all $i$, then $|S|\leq 2k$. Note that $|S\cap V_{i}|=4$ is impossible because $V_{i}=N[b_{i}]\not\subseteq S$ by Lemma 2.9. We show that if $|S\cap V_{i}|=3$, then $|S\cap V_{i-1}|,|S\cap V_{i+1}|\leq 1$. Once this is established, we conclude that $|S|\leq 2k$.
First, we assume that $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j}\cap S$, and show $|V_{j}\cap S|\leq 1$. Let $x\in\{b_{j},c_{j},d_{j}\}\cap S$ and let $F_{x}$ be a private fort of $x$. If $F_{x}\cap\{b_{j},d_{j}\}\neq\varnothing$, then $a_{j}$ must be adjacent to two vertices in the fort and $c_{j-1}\not\in F_{x}$, so $\{b_{j},d_{j}\}\subseteq F_{x}$. If $x=c_{j}$, then $b_{j}\in F_{x}$ or $b_{j}$ needs two neighbors in $F_{x}$ so $d_{j}\in F_{x}$, so $\{b_{j},c_{j},d_{j}\}\subseteq F_{x}$. Therefore $|V_{j}\cap S|\leq 1$. By symmetry, if $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j-1}\cap S$, then $|S\cap V_{j-1}|\leq 1$. Thus {linenomath}
$c_{j-1},a_{j}\mbox{ cannot be in a private fort of any }v\in V_{j}\cap S$ $\implies|V_{j}\cap S|\leq 1.$ (4.1)
$c_{j-1},a_{j}\mbox{ cannot be in a private fort of any }v\in V_{j-1}\cap S$ $\implies|V_{j-1}\cap S|\leq 1.$ (4.2)
Now suppose that $|V_{i}\cap S|=3$. First assume that $a_{i},b_{i},c_{i}\in S$. Note that $d_{i}$ cannot belong to a private fort of a vertex not in $V_{i}$, since then $b_{i}$ would be adjacent to only one vertex in the fort. Because $d_{i}$ (and $b_{i}$) cannot be in a private fort of a vertex not in $V_{i}$, $c_{i-1}$ cannot belong to any private fort of a vertex in outside $V_{i}$ (or $a_{i}$ would have only one neighbor in this fort). Thus $a_{i}$ and $c_{i-1}$ cannot belong to any private fort of a vertex in $V_{i-1}$, and so $|S\cap V_{i-1}|\leq 1$. By symmetry, $|S\cap V_{i+1}|\leq 1$.
Now assume that $b_{i},c_{i},d_{i}\in S$. Note that $a_{i}$ cannot belong to a private fort of any vertex outside $V_{i}$, since then $b_{i}$ and $d_{i}$ would be adjacent to only one vertex in that fort. Furthermore, if $c_{i-1}$ is in a private fort of a vertex not in $V_{i}$, $a_{i}$ would only be adjacent to one vertex in that fort. Similarly, $a_{i+1}$ cannot belong to a private fort of a vertex not in $V_{i}$, since then $c_{i}$ would be adjacent to only one vertex in that fort. By (4.1) and (4.2), $|S\cap V_{i-1}|\leq 1$ and $|S\cap V_{i+1}|\leq 1$. ∎
Remark 4.5.
It is known that $\operatorname{Z}(N_{k})=k+2$ [15, Theorem 9.5]. Suppose $S$ is a minimal zero forcing set. Since $b_{i}$ and $d_{i}$ are twins, at least one of $b_{i}$ and $d_{i}$ must be in $S$ by Proposition 1.5. In order for the forcing to start, there must be a blue vertex with three blue neighbors so there must be at least two other vertices in $S$, and with these vertices we have a zero forcing set. Thus $\overline{\operatorname{Z}}(N_{k})=k+2$.
Theorem 4.6.
[10] Let $G$ be a graph of order $n$. If $\delta(G)=2$, then $\gamma_{2}(G)\leq\frac{2n}{3}$.
Corollary 4.7.
Let $G$ be a graph of order $n\geq 3$ and $\delta(G)=2$. Then $\operatorname{ZIR}(G)>\frac{n}{3}$.
Proof.
From Proposition 4.1 and Theorem 4.6 it follows that $\operatorname{ZIR}(G)\geq\frac{n}{3}$. Graphs with 2-domination number equal to $\frac{2n}{3}$ were characterized in [14] as $G=H\circ K_{2}$ for any graph $H$. We show that $\operatorname{ZIR}(G)>\frac{n}{3}$.
Label the vertices of $H$ as $v_{1},v_{2},\dots,v_{n_{H}}$ and the vertices of $K_{2}$ joined to vertex $v_{i}$ as $a_{i}$ and $b_{i}$. The set $S=\{v_{1}\}\cup\{a_{i}:i=1,\dots,n_{H}\}$ is a ZIr-set of $G$ where the private fort of each $a_{i}$ is $\{a_{i},b_{i}\}$ and the private fort of $v_{1}$ is $V(G)\setminus\{a_{i}:i=1,\dots,n_{H}\}$. ∎
Theorem 5.11 provides a family of graphs $G$ satisfying $\operatorname{ZIR}(G)=\frac{1}{3}|V(G)|$, but these graphs have $\delta(G)=1$, so we do not know how close to the bound $\frac{1}{3}|V(G)|$ we can get for graphs $G$ with $\delta(G)=2$.
The next family of graphs realizes $\operatorname{ZIR}(G)=n-\gamma_{2}(G)$ and $\delta(G)=2$. Let $H_{k}$ be the graph consisting of $k$ 5-cycles, where the vertices of the $i$-th cycle is labeled $v_{i,1},v_{i,2},\dots,v_{i,5}$, together with the edges $v_{i,3}v_{i+1,1}$ for $i\in\{1,\dots,k-1\}$ and the edge $v_{k,3}v_{1,1}$; see Figure 4.2.
$v_{1,3}$$v_{1,2}$$v_{1,1}$$v_{1,5}$$v_{1,4}$$v_{2,3}$$v_{2,2}$$v_{2,1}$$v_{2,5}$$v_{2,4}$$v_{3,3}$$v_{3,2}$$v_{3,1}$$v_{3,5}$$v_{3,4}$
Figure 4.2: The graph $H_{3}$
Theorem 4.8.
The graph $H_{k}$ of order $n=5k$ has $\delta(H_{k})=2$, $\gamma_{2}(H_{k})=\frac{3n}{5}$, and $\operatorname{ZIR}(H_{k})=\frac{2n}{5}=n-\gamma_{2}(H_{k})$. Furthermore, $\operatorname{Z}(H_{k})=k+2$.
Proof.
Let $B_{i}$ denote the set of vertices of the $i$-th cycle of $H_{k}$. First we show that $\gamma_{2}(H_{k})=\frac{3n}{5}$. Let $D$ be a 2-dominating set of $H_{k}$. If $v_{i,2}\in D$, then to 2-dominate $v_{i,4}$ and $v_{i,5}$ $|D\cap B_{i}\setminus\{v_{i,2}\}|\geq 2$. It therefore follows that $|D\cap B_{i}|\geq 3$. On the other hand, if $v_{i,2}\not\in D$, then certainly $v_{i,1},v_{i,3}\in D$. Since $v_{i,4}$ and $v_{i,5}$ are not 2-dominated yet, it follows that $|D\cap\{v_{i,4},v_{i,5}\}|\geq 1$ and so $|D\cap B_{i}|\geq 3$. Hence $\gamma_{2}(H_{k})\geq\frac{3n}{5}$. Since $\{v_{i,2},v_{i,4},v_{i,5}:i=1,\dots,k\}$ is a 2-dominating set, equality follows.
Since $\operatorname{ZIR}(H_{k})\geq\frac{2n}{5}$, it suffices to show that $\operatorname{ZIR}(H_{k})\leq\frac{2n}{5}$. Let $S$ be a ZIr-set of $H_{k}$. If $|S\cap B_{i}|\leq 2$ for every $i\in\{1,\dots,k\}$ then the result follows. Assume therefore that $|S\cap B_{i}|\geq 3$ for some $i\in\{1,\dots,k\}$. If $|S\cap B_{i}|>3$, there exists a degree-2 vertex $v$ in $B_{i}$ such that $N[v]\subseteq S$, which is impossible according to Lemma 2.9. Therefore $|S\cap B_{i}|=3$. To complete the proof, we show that $|S\cap B_{i-1}|\leq 1$ and $|S\cap B_{i+1}|\leq 1$. Let $F_{i,j}$ denote a private fort of $v_{i,j}$ (relative to $S$).
First we examine $B_{i}\cap S$ further, showing $v_{i,1}$ and $v_{i,3}$ cannot both be in $S$. To obtain a contradiction, assume $\{v_{i,1},v_{i,3}\}\subseteq S$. Since $N(v_{i,2})=\{v_{i,1},v_{i,3}\}$, $v_{i,2}\in F_{i,1}$. If $v_{i,5}\in F_{i,1}$, then $v_{i,4}$ would necessarily also be in $F_{i,1}$, contradicting $|S\cap B_{i}|=3$. So $v_{i,5}\not\in F_{i,1}$, and it follows that $v_{i,4}\in F_{i,1}$ to ensure that $v_{i,5}$ is adjacent to at least two vertices in $F_{i,1}$. A similar argument for a private fort $F_{i,3}$ shows that $v_{i,5}\in F_{i,3}$. In this case, $B_{i}\subseteq F_{i,1}\cup F_{i,3}$. Thus $|B_{i}\cap S|=2$, which is a contradiction. Hence |$|\{v_{i,1},v_{i,3}\}\cap S|\leq 1$.
For the two cases |$|\{v_{i,1},v_{i,3}\}\cap S|=0$ and |$|\{v_{i,1},v_{i,3}\}\cap S|=1$, we prove both the following statements:
  1. $(a)$
    There are $j$ and $j^{\prime}$ such that $v_{i+1,1}\in F_{i,j}$, and $v_{i-1,3}\in F_{i,j^{\prime}}$.
  2. $(b)$
    $v_{i+1,1}\not\in F_{i+1,j}$ for any $j$ and $v_{i-1,3}\not\in F_{i-1,j}$ for any $j$.
Case |$|\{v_{i,1},v_{i,3}\}\cap S|=0$:
First, we prove (a) for both $v_{i+1,1}$ and $v_{i-1,3}$. Since $|B_{i}\cap S|=3$, $B_{i}\cap S=\{v_{i,2},v_{i,4},v_{i,5}\}$. Then $v_{i,3}$ is the only neighbor of $v_{i,4}$ not in $S$ and therefore $v_{i,3}\not\in F_{i,2}$. To ensure that $v_{i,3}$ is adjacent to two vertices of $F_{i,2}$, $v_{i+1,1}\in F_{i,2}$. Similarly, $v_{i-1,3}\in F_{i,2}$.
To prove (b), suppose $v_{i+1,1}\in F_{i+1,j}$ for some $j\in\{1,\dots,5\}$. Since both of the neighbors of $v_{i,3}$ in $B_{i}$ are in $S$, $v_{i,3}$ is necessarily in $F_{i+1,j}$ but then $v_{i,4}$ is not adjacent to two vertices of $F_{i+1,j}$. Therefore $v_{i+1,1}\not\in F_{i+1,j}$ for any $j$ and similarly $v_{i-1,3}\not\in F_{i-1,j}$ for any $j$.
Case |$|\{v_{i,1},v_{i,3}\}\cap S|=1$:
Assume that $\{v_{i,1},v_{i,3}\}\cap S=\{v_{i,3}\}$. It follows from Lemma 2.9 that $N[v_{i,4}]\not\subseteq S$ and thus at most one of $v_{i,4}$ or $v_{i,5}$ is in $S$. Since $|B_{i}\cap S|=3$, $v_{i,2}\in S$ and exactly one of $v_{i,4}$ or $v_{i,5}$ is in $S$.
First, we prove (a) for $v_{i+1,1}$. Either $v_{i,4}\in S$ or $v_{i,4}\in F_{i,5}$. In both cases $v_{i,3}$ is adjacent to a vertex in $F_{i,j}$ where $j=4$ or $5$, and therefore $v_{i+1,1}\in F_{i,j}$.
Next, we prove (a) for $v_{i-1,3}$. Let $j^{\prime}=4$ or $5$, according to which of $v_{i,4}$ or $v_{i,5}$ is in $S$. If $v_{i,1}\in F_{i,j^{\prime}}$, then $v_{i,2}$ would have only one neighbor in $F_{i,j^{\prime}}$. Thus $v_{i,1}\not\in F_{i,j^{\prime}}$, and $v_{i-1,3}\in F_{i,j^{\prime}}$ to ensure that $v_{i,1}$ is adjacent to two vertices of $F_{i,j^{\prime}}$.
To prove (b), suppose $v_{i+1,1}\in F_{i+1,j}$ for some $j\in\{1,\dots,5\}$. Since $v_{i,2},v_{i,3}\in S$, $v_{i,4}$ is necessarily in $F_{i+1,j}$ and therefore $v_{i,5}\in S$. To ensure that $v_{i,5}$ is adjacent to two vertices of $F_{i+1,j}$, $v_{i,1}\in F_{i+1,j}$ but then $v_{i,2}$ is only adjacent to one vertex of $F_{i+1,j}$. Therefore $v_{i+1,1}\not\in F_{i+1,j}$ for any $j$. Now suppose that $v_{i-1,3}\in F_{i-1,j}$ for some $j\in\{1,\dots,5\}$. Since $v_{i,2},v_{i,3}\in S$, $v_{i,1}\not\in F_{i-1,j}$ and therefore $v_{i,5}\in F_{i-1,j}$. But then $v_{i,4}\in S$ and only adjacent to one vertex of $F_{i-1,j}$. Hence $v_{i-1,3}\not\in F_{i-1,j}$ for any $j$.
To complete the proof we show that $|S\cap B_{i+1}|\leq 1$ and $|S\cap B_{i-1}|\leq 1$.
From (a), $v_{i+1,1}\in F_{i,\ell}$ for some specific $\ell\in\{1,\dots,5\}$ and it therefore follows that |$|\{v_{i+1,2},v_{i+1,3}\}\cap F_{i,\ell}|\geq 1$ and |$|\{v_{i+1,4},v_{i+1,5}\}\cap F_{i,\ell}|\geq 1$.
Now consider the vertices in $S\cap B_{i+1}$. From (b) it follows that $v_{i+1,1}\not\in F_{i+1,j}$ for any $j\in\{1,\dots,5\}$, which implies $v_{i+1,1}\not\in S$. Also we can see that $v_{i,3}\not\in F_{i+1,j}$ either: If $v_{i,3}\in S$ then clearly $v_{i,3}\not\in F_{i+1,j}$. So assume $v_{i,3}\not\in S$ and $v_{i,3}\in F_{i+1,j}$. If $v_{i,1}\not\in S$, then $v_{i,2},v_{i,4},v_{i,5}\in S$ and $v_{i,4}$ is adjacent to only one vertex in $F_{i+1,j}$. If $v_{i,1}\in S$, then also $v_{i,2}\in S$ and $v_{i,2}$ is adjacent to only one vertex in $F_{i+1,j}$. Thus $v_{i,3}\not\in F_{i+1,j}$.
Since $v_{i,3}\not\in F_{i+1,j}$ for any $j$, $\{v_{i+1,2},v_{i+1,5}\}\not\subseteq S$, otherwise $v_{i+1,1}$ is adjacent to only one vertex in $F_{i+1,2}$ and $F_{i+1,5}$.
Suppose first that $\{v_{i+1,2},v_{i+1,5}\}\cap S=\{v_{i+1,2}\}$. Then $v_{i+1,5}\in F_{i+1,2}$ to ensure that $v_{i+1,1}$ is adjacent to two vertices of $F_{i+1,2}$. Recall $v_{i+1,1}\in F_{i,\ell}$, so $v_{i+1,3}\in F_{i,\ell}$ to ensure that $v_{i+1,2}$ is adjacent to two vertices of $F_{i,\ell}$. If also $v_{i+1,4}\in S$, then $v_{i+1,5}\in F_{i+1,4}$. This is impossible since $v_{i+1,1}$ is then adjacent to only one vertex of $F_{i+1,4}$. Thus $\{v_{i+1,2},v_{i+1,5}\}\cap S=\{v_{i+1,2}\}$ implies $|S\cap B_{i+1}|\leq 1$. If $\{v_{i+1,2},v_{i+1,5}\}\cap S=\{v_{i+1,5}\}$ it follows similarly that $|S\cap B_{i+1}|\leq 1$.
Now, if $\{v_{i+1,2},v_{i+1,5}\}\cap S=\emptyset$, then $|S\cap B_{i+1}|\leq 2$. Assume $|S\cap B_{i+1}|=2$ with $v_{i+1,3},v_{i+1,4}\in S$. Then $v_{i+1,2},v_{i+1,5}\in F_{i,\ell}$ since $v_{i+1,1}\in F_{i,\ell}$ and $v_{i+1,3},v_{i+1,4}\not\in F_{i,\ell}$. But then $v_{i+1,4}$ is adjacent to only one vertex in $F_{i,\ell}$, a contradiction. It follows that $|S\cap B_{i+1}|\leq 1$. It follows similarly that $|S\cap B_{i-1}|\leq 1$ and hence $\operatorname{ZIR}(H_{k})\geq\frac{2}{5}n$ as required.
For the zero forcing number, observe that $\{v_{1,3},v_{1.4}\}\cup\{v_{i,2}:i=1,\dots,k\}$ is a zero forcing set of size $k+2$. Since $\{v_{i,2},v_{i,4},v_{i,5}\}$ is a fort, a zero forcing set must contain at least one vertex of each cycle. Since $\deg(v_{i,1})=\deg(v_{i,3})=3$, forcing cannot move from one cycle to the next unless some cycle has three vertices or two adjacent cycles have two each. Thus $\operatorname{Z}(H_{k})=k+2$. ∎
Finally, we present bounds that do not involve $\gamma_{2}(G)$.
Remark 4.9.
Note that $\operatorname{Z}(G)\leq\operatorname{ZIR}(G)$, and it is known that $\operatorname{Z}(G(n,p))=n-o(n)$ where $G(n,p)$ is the Erdős-Rényi random graph of order $n$ with edge probability $p$ [15, Section 9.11].
Proposition 4.10.
Let $G$ be a connected graph of order $n\geq 2$. Then $\operatorname{ZIR}(G)\leq\frac{\Delta(G)}{\Delta(G)+1}n$ and this bound is sharp.
Proof.
Let $S$ be a maximal ZIr-set of $G$. Let $X$ denote the set of edges incident to a vertex in $S$ and a vertex in $V\setminus S$. Since $G$ is connected, a vertex $x\in S$ has a neighbor $w$. If $w\in S$, then $x$ has another neighbor $u\in F_{w}$ (the private fort of $w$), and $u\not\in S$. Thus every vertex in $S$ is adjacent to at least one vertex not in $S$ and therefore $|X|\geq|S|$. Furthermore, every vertex in $V\setminus S$ is adjacent to at most $\Delta(G)$ vertices of $S$ and therefore $\Delta(G)|V\setminus S|\geq|X|$. It follows that
$\Delta(G)(n-|S|)=\Delta(G)|V\setminus S|\geq|X|\geq|S|.$
Since $S$ is any maximal ZIr-set, $\operatorname{ZIR}(G)\leq\frac{\Delta(G)}{\Delta(G)+1}n$. By Example 2.12, $\operatorname{ZIR}(K_{n})=n-1=\frac{n-1}{(n-1)+1}n$. ∎
Corollary 4.11.
If $G$ is a connected cubic graph of order $n\geq 4$, then $\frac{1}{2}n\leq\operatorname{ZIR}(G)\leq\frac{3}{4}n$ and the lower bound is sharp.
Proof.
The inequalities are immediate from Corollary 4.3 and Proposition 4.10. The lower bound is sharp by Theorem 4.4. ∎
The only example we know where the upper bound is equality is $\operatorname{ZIR}(K_{4})=3$.
5 Extreme values of $\operatorname{zir}$ and $\operatorname{ZIR}$
In this section we characterize graphs of order $n$ having the extreme high values $n$ and $n-1$ for $\operatorname{zir}$ and $\operatorname{ZIR}$ and characterize graphs having the extreme high value $n-2$ for $\operatorname{zir}$. We also characterize graphs having the extreme low value 1 for $\operatorname{zir}$ and present a family of graphs with the lowest known $\operatorname{ZIR}$ value for arbitrarily large $n$.
Remark 5.1.
Let $G$ be a graph of order $n$. It is well known that $\operatorname{Z}(G)=n\Leftrightarrow G\cong K_{n}\Leftrightarrow\overline{\operatorname{Z}}(G)=n$, and this extends to $\operatorname{zir}$ and $\operatorname{ZIR}$: Since $\operatorname{zir}(G)=n$ implies $\operatorname{Z}(G)=n$ and $\operatorname{ZIR}(G)=n$ implies $G\cong\overline{K_{n}}$ by Remark 2.7, we have $G\cong\overline{K_{n}}\Leftrightarrow\operatorname{zir}(G)=n\Leftrightarrow\operatorname{ZIR}(G)=n$.
For a graph on $n$ vertices, it is well known that $\operatorname{Z}(G)=n-1$ if and only if $G\cong K_{n-r}\mathbin{\,\sqcup\,}rK_{1}$ with $n-r\geq 2$. It was shown recently by Brimkov and Carlson in [7] that $\overline{\operatorname{Z}}(G)=n-1$ if and only if $G\cong K_{n-r}\mathbin{\,\sqcup\,}rK_{1}$. We extend this to $\operatorname{zir}$ and $\operatorname{ZIR}$.
Proposition 5.2.
Let $G$ be a graph of order $n\geq 2$. The following are equivalent:
  1. 1.
    $G\cong K_{n-r}\mathbin{\,\sqcup\,}rK_{1}$ with $n-r\geq 2$.
  2. 2.
    $\operatorname{zir}(G)=n-1$.
  3. 3.
    $\operatorname{Z}(G)=n-1$.
  4. 4.
    $\overline{\operatorname{Z}}(G)=n-1$.
  5. 5.
    $\operatorname{ZIR}(G)=n-1$.
Proof.
If $G$ does not have an edge, then $G\cong\overline{K_{n}}$ and $\operatorname{zir}(\overline{K_{n}})=\operatorname{Z}(\overline{K_{n}})=\overline{\operatorname{Z}}(\overline{K_{n}})=\operatorname{ZIR}(\overline{K_{n}})=n$. So assume $G$ has an edge, so that $\operatorname{ZIR}(G)\leq n-1$ by Remark 2.7. Then with Example 2.12 and Corollary 2.3, we have that (1) $\Rightarrow$ (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5). Since $\overline{\operatorname{Z}}(H)=|V(H)|-1$ implies $H\cong K_{n-r}\mathbin{\,\sqcup\,}rK_{1}$ with $n-r\geq 2$ for any graph $H$ of order $n$ [7], (4) $\Rightarrow$ (1). Thus (1) – (4) are equivalent.
So assume $\operatorname{ZIR}(G)=n-1$. Then there is a ZIr-set $S$ such that $|S|=n-1$. Since every set of $n-1$ vertices is a zero forcing set of $G$, $S$ is a minimal zero forcing set by Remark 2.1. Thus $\overline{\operatorname{Z}}(G)=n-1$. ∎
The next result provides many examples of graphs with $\operatorname{ZIR}(G)=|V(G)|-2$.
Corollary 5.3.
If $G=H\vee 2K_{1}$ or $G=H\vee{K_{2}}$ with $H\not\cong K_{|V(H)|}$, then $\operatorname{ZIR}(G)=|V(H)|=|V(G)|-2$. If $G=H\vee{K_{2}}$ and $H$ has no isolated vertices, then $\operatorname{Z}(G)=\operatorname{Z}(H)+2$ and $\overline{\operatorname{Z}}(G)=\overline{\operatorname{Z}}(H)+2$.
Proof.
The statement about $\operatorname{ZIR}$ is immediate from the previous proposition and Proposition 4.1 (because $V(G)\setminus V(H)$ is a 2-dominating set). It is known that $\operatorname{Z}(H\vee K_{1})=\operatorname{Z}(H)+1$ for a graph $H$ with no isolated vertices [15, Proposition 9.16]. Since $H\vee K_{2}\cong(H\vee K_{1})\vee K_{1}$, this implies $\operatorname{Z}(H\vee K_{2})=\operatorname{Z}(H)+2$ if $H$ has no isolated vertices. Furthermore, the proof of [15, Proposition 9.16] shows that $\overline{\operatorname{Z}}(H\vee K_{1})=\overline{\operatorname{Z}}(H)+1$ for a graph $H$ with no isolated vertices, which implies $\overline{\operatorname{Z}}(G)=\overline{\operatorname{Z}}(H)+2$ for $G=H\vee K_{2}$. ∎
The previous corollary does not list $\operatorname{Z}(H\vee 2K_{1})$ because $\operatorname{Z}(H)+1\leq\operatorname{Z}(H\vee 2K_{1})\leq Z(H)+2$ and both bounds are sharp (consider $H\cong K_{r}$ for the lower bound and $H\cong P_{r}$ for the upper bound). It is not true that all graphs $G$ of order $n$ with $\operatorname{ZIR}(G)=n-2$ have $\gamma_{2}(G)=2$, as seen in the next example.
Example 5.4.
Denote the vertices of the partite sets of $K_{3,4}$ by $U=\{u_{1},u_{2},u_{3}\}$ and $\{x_{1},x_{2},y_{1},y_{2}\}$. Construct $G$ from $K_{3,4}$ by adding the edges $x_{1}y_{1}$ and $x_{2}y_{2}$. See Figure 5.1. Note that no set of two vertices 2-dominates $G$ since the degree of every vertex is four. Since $U$ is a 2-dominating set, $\gamma_{2}(G)=3$. We see that $S=\{u_{1},u_{2},u_{3},y_{1},y_{2}\}$ is a ZIr-set by noting that $F_{u_{i}}=\{u_{i},x_{1},x_{2}\}$ is a private fort of $u_{i}$ for $i=1,2,3$ and $F_{y_{j}}=\{y_{j},x_{j}\}$ is a private fort of $y_{j}$ for $j=1,2$. Thus $\operatorname{ZIR}(G)=5=|V(G)|-2$.
$u_{1}$$u_{2}$$u_{3}$$x_{1}$$y_{1}$$x_{2}$$y_{2}$
Figure 5.1: A graph $G$ with $\operatorname{ZIR}(G)=|V(G)|-2$ and $\gamma_{2}(G)=3$.
There is a known characterization of $\operatorname{Z}(G)\geq n-2$ for graphs $G$ of order $n\geq 3$. In [2] it is shown that $\operatorname{Z}(G)\geq n-2$ if and only if $G$ does not contain $P_{4}$, $P_{3}\mathbin{\,\sqcup\,}K_{2}$, ${\rm dart}$, $\ltimes$, or $3K_{2}$ as an induced subgraph. In [4] it is shown that $G$ does not contain $P_{4}$, $P_{3}\mathbin{\,\sqcup\,}K_{2}$, ${\rm dart}$, $\ltimes$, or $3K_{2}$ as an induced subgraph if and only if
$G\cong\overline{\left(K_{s_{1}}\mathbin{\,\sqcup\,}\dots\mathbin{\,\sqcup\,}K_{s_{t}}\mathbin{\,\sqcup\,}K_{q_{1},p_{1}}\mathbin{\,\sqcup\,}K_{q_{2},p_{2}}\mathbin{\,\sqcup\,}\cdots\mathbin{\,\sqcup\,}K_{q_{k},p_{k}}\right)\vee K_{r}}$
with $s_{i}\geq 1$, $p_{i}\geq 1$, $q_{i}\geq 0$, $t,k,r\geq 0$, and $t+k+r\geq 1$. Observe that $K_{1}\cong K_{0,1}$ and $K_{2}\cong K_{1,1}$, so we may assume $s_{i}\geq 3$. Without loss of generality we may also assume $p_{i}\geq q_{i}$ (which implies $p_{i}\geq 1$) and $q_{i}\geq q_{i+1}$. Furthermore, $K_{0,a_{1}}\mathbin{\,\sqcup\,}\dots\mathbin{\,\sqcup\,}K_{0,a_{\ell}}\cong K_{0,a_{1}+\dots+a_{\ell}}$, so we may assume $q_{k-1}\geq 1$. These results can be combined as in the next theorem.
Theorem 5.5.
[2, 4] For a graph $G$ of order $n\geq 3$, $\operatorname{Z}(G)\geq n-2$ if and only if
$\overline{G}\cong\left(K_{s_{1}}\mathbin{\,\sqcup\,}\dots K_{s_{t}}\mathbin{\,\sqcup\,}K_{q_{1},p_{1}}\mathbin{\,\sqcup\,}K_{q_{2},p_{2}}\mathbin{\,\sqcup\,}\cdots\mathbin{\,\sqcup\,}K_{q_{k},p_{k}}\right)\vee K_{r}$
with $s_{i}\geq 3$, $p_{i}\geq q_{i}$, $q_{i}\geq q_{i+1}$, $p_{i}\geq 1$, $q_{k-1}\geq 1$ and $q_{k}\geq 0$, $t,k,r\geq 0$, and $t+k+r\geq 1$.
Note that $\operatorname{Z}(G)\geq|V(G)|-2$ implies $\operatorname{ZIR}(G)\geq|V(G)|-2$. For example, the graph $G$ in Example 5.4 is covered by Theorem 5.5 since $\overline{G}=K_{3}\mathbin{\,\sqcup\,}K_{2,2}$. However, $\operatorname{ZIR}(G)\geq|V(G)|-2$ does not imply $\operatorname{Z}(G)\geq|V(G)|-2$, as seen in the next example.
Example 5.6.
Consider the graph $G=rK_{2}\vee 2K_{1}$ with $r\geq 3$. Corollary 5.3 establishes $\operatorname{ZIR}(G)=2r=|V(G)|-2$. To see that $\operatorname{Z}(G)\leq r+2$, note that a set consisting of one vertex from each $K_{2}$ and the vertices $u$ and $u^{\prime}$ of $2K_{1}$ is a zero forcing set of cardinality $r+2$, so $\operatorname{Z}(G)\leq r+2<2r=|V(G)|-2$. Furthermore, $\operatorname{Z}(G)=r+2$, because any zero forcing set must contain at least one vertex of each $K_{2}$ and at least one of $u,u^{\prime}$, and one more vertex (either contain both $u$ and $u^{\prime}$ or contain both vertices of some $K_{2}$).
We use Theorem 5.5 to characterize graphs $G$ having $\operatorname{zir}(G)=|V(G)|-2$.
Theorem 5.7.
Let $G$ be a graph of order $n\geq 3$. Then $\operatorname{zir}(G)=n-2$ if and only if
$\overline{G}\cong{\left(K_{q_{1},p_{1}}\mathbin{\,\sqcup\,}K_{q_{2},p_{2}}\mathbin{\,\sqcup\,}\cdots\mathbin{\,\sqcup\,}K_{q_{k},p_{k}}\right)\vee K_{r}}$
with $p_{i}\geq q_{i}$, $q_{i}\geq q_{i+1}$, $p_{i}\geq 1$, $q_{k-1}\geq 1$ and $q_{k}\geq 0$, $k\geq 1$, $r\geq 0$, and
$q_{1}\geq 2\ \mbox{ or }\ (q_{1}=1\mbox{ and }k\geq 2).$
Proof.
Note that $\operatorname{zir}(G)=n-2$ implies $\operatorname{Z}(G)\geq n-2$. Thus it suffices to assume $G$ has the form in Theorem 5.5, and determine which of these forms have $\operatorname{zir}(G)=n-2$. Recall $\operatorname{zir}(G)=n$ if and only if $G\cong nK_{1}$, and $\operatorname{zir}(G)=n-1$ if and only if $G\cong K_{n-r}\mathbin{\,\sqcup\,}rK_{1}$ with $n-r\geq 2$. This eliminates several cases: $t=1,k=0$ means $G\cong(s_{1}+r)K_{1}$; $t=0,k=1,q_{1}=0$ means $G\cong K_{p_{1}}\mathbin{\,\sqcup\,}rK_{1}$; $t=0,k=1,q_{1}=1$ means $G\cong K_{p_{1}}\mathbin{\,\sqcup\,}(r+1)K_{1}$. Thus $t=1,k=0$ and $t=0,k=1,q_{1}=0,1$ imply $\operatorname{zir}(G)>n-2$. This means $G$ has one of the following forms:
  1. (a)
    $t\geq 2$.
  2. (b)
    $t=1$ and $k\geq 1$.
  3. (c)
    $t=0$ and $q_{1}\geq 2$.
  4. (d)
    $t=0$, $k\geq 2$, and $q_{1}=1$.
Observe that the presence or absence of $r$ isolated vertices in $G$ ($\vee K_{r}$ in $\overline{G}$) has no effect on whether or not $\operatorname{zir}(G)=n-2$, because each isolated vertex is its own private fort. Thus we assume $r=0$.
We first assume we have one of cases (a) and (b) and show that $\operatorname{zir}(G)<n-2$. In both these cases, $t\geq 1$ and $s_{1}\geq 3$. Let $U$ denote the vertices of $\overline{K_{s_{1}}}$. Since $t\geq 2$ or $k\geq 1$, $G$ is connected. Let $S=V(G)\setminus U$; note $S\neq\emptyset$. Since $G[U]$ is the empty graph on $s_{1}$ vertices and every vertex in $U$ is adjacent to every vertex not in $U$, $U$ is in every private fort of a vertex in $S$. Therefore $S$ is a maximal ZIr-set, where $\{v\}\cup U$ is a private fort of $v\in S$ with $|S|=n-|U|<n-2$.
Case (c) and (d) satisfy the description of $G$ in the statement of the theorem, so it remains to show that every graph $G$ of this form satisfies $\operatorname{zir}(G)=n-2$. Denote the partite sets as $Z_{1},\dots,Z_{2k}$, or $Z_{2k-1}$ if $q_{k}=0$, where $|Z_{2j-1}|=p_{j}$ and $|Z_{2j}|=q_{j}$ for $j=1,\dots,k$. Let $S$ be a maximal ZIr-set. We show that $|S|\geq n-2$, which implies $\operatorname{zir}(G)=n-2$. The only case in which a partite set is a dominating set is if $q_{k}=0$; in this case each vertex of $Z_{2k-1}$ dominates $G$. Suppose first that $Z_{2k-1}$ is a dominating set and $\left(V(G)\setminus S\right)\subseteq Z_{2k-1}$. If $|V(G)\setminus S|\leq 2$, then $|S|\geq n-2$. If $|V(G)\setminus S|\geq 3$ then $S$ would not be maximal, because for any distinct $u,w\in Z_{2k-1}$, $V\setminus\{u,w\}$ is a ZIr-set since $\{v,u,w\}$ is a private fort of $v$. So assume $\left(V(G)\setminus S\right)\not\subseteq Z_{2k-1}$ or $Z_{2k-1}$ is not a dominating set. For all $i$, if $Z_{i}$ is not a dominating set of $G$, then $\left(V(G)\setminus S\right)\not\subseteq Z_{i}$ by Lemma 2.9. This means $\left(V(G)\setminus S\right)\not\subseteq Z_{i}$ for all $i$, so there exist distinct indices $j$ and $\ell$ and vertices $z_{j}\in\left(V(G)\setminus S\right)\cap Z_{j}$ and $z_{\ell}\in\left(V(G)\setminus S\right)\cap Z_{\ell}$. We show that $V\setminus\{z_{j},z_{\ell}\}$ is a ZIr-set, so the maximality of $S$ implies $V\setminus\{z_{j},z_{\ell}\}\subseteq S$ and $|S|\geq n-2$. Let $v\neq z_{j},z_{\ell}$ and $v\in Z_{i}$. If $i\neq j,\ell$, the $\{v,z_{j},z_{\ell}\}$ is a private fort of $v$. If $i=j$, then $\{v,z_{j}\}$ is a private fort of $v$, and the case $i=\ell$ is similar. ∎
Remark 5.8.
Suppose $\operatorname{zir}(G)=1$ and $\{x\}$ is a maximal ZIr-set. If $F_{x}$ is a fort that contains $x$, then $F\subseteq F_{x}$ for every fort $F$ that does not contain $x$, because if $F\not\subseteq F_{x}$, then there exists $y\in F\setminus\ F_{x}$ and $\{x,y\}$ is a ZIr-set.
Theorem 5.9.
Let $G$ be a graph of order $n\geq 1$. Then $\operatorname{zir}(G)=1$ if and only if $G=P_{n}$ or $G=K_{1,n-1}$.
Proof.
If $G=P_{n}$ or $G=K_{1,n-1}$, then $\operatorname{zir}(G)=1$ by Proposition 3.3 or Example 3.1. Assume $\operatorname{zir}(G)=1$. If $G$ had more than one connected component, then $\operatorname{zir}(G)\geq 2$ by Remark 2.7, so $G$ is connected. If $n=1$ then $G=P_{1}$. So assume $n\geq 2$ and $G\neq K_{1,p}$ with $p\geq 1$. Then $G$ has at least two vertices of degree at least two.
Since $\operatorname{zir}(G)=1$, there exists a maximal ZIr-set $\{x\}$. By Proposition 2.10, $\deg x\not\geq 2$ since $\{x\}$ is maximal and there are two vertices of degree two or more. Thus $\deg x=1$.
Now suppose that $G$ has a vertex of degree at least three. Let $c$ be a vertex such that $\deg c\geq 3$ and $\deg w\geq 3$ implies $\operatorname{dist}(x,w)\geq\operatorname{dist}(x,c)$. Label the vertices on the unique path from $c$ to $x$ by $v_{1},\dots,v_{k}$ with $\operatorname{dist}(v_{i},c)=i$. Since $c$ is a cut-vertex and $\deg c\geq 3$, $F_{1}=V(G)\setminus\{c,v_{1},\dots,v_{k},x\}$ is a fort. Since $x\not\in F_{1}$ and $\{x\}$ is a maximal ZIr-set, it follows from Remark 5.8 that every fort that contains $x$ must also contain $F_{1}$. For any vertex $u$, $\deg u\geq 2$ implies $F_{2}=V(G)\setminus\{u\}$ is a fort that contains $x$. If $u\neq v_{i}$, then $F_{1}\not\subseteq F_{2}$. Thus $\deg u\geq 2$ implies $u=v_{i}$ for some $i$. Since $\deg c\geq 3$, there are two vertices $w_{1},w_{2}\in N(c)$ with $w_{1},w_{2}\neq v_{1}$. Thus $\deg w_{1}=\deg w_{2}=1$. Thus $F_{3}=V(G)\setminus\{c,w_{1}\}$ is a fort that contains $x$. But $F_{1}\not\subseteq F_{3}$, which is a contradiction. Thus $G$ cannot have a vertex of degree greater than two. Since $\deg x=1$, $G$ is a path. ∎
Note that $\operatorname{ZIR}(G)=1$ implies $\operatorname{Z}(G)=1$ and thus $G\cong P_{n}$. But $\operatorname{ZIR}(P_{n})=1$ only when $n\leq 3$ (see Proposition 3.3 and Observation 3.4).
We do not know what the lowest value of $\operatorname{ZIR}(G)$ is over all graphs of order $n$ as $n\to\infty$. A computer search of graphs of order $n\leq 9$ found several graphs $G$ with order 9 and $\operatorname{ZIR}(G)=3$, including $C_{3}\circ 2K_{1}$. This led to Theorem 5.11 below that $\operatorname{ZIR}(C_{r}\circ 2K_{1})=\frac{1}{3}|V(C_{r}\circ 2K_{1})|$, the lowest known value of $\operatorname{ZIR}(G)$ among graphs $G$ of order $n$ as $n$ becomes arbitrarily large. Before stating Theorem 5.11, we introduce some notation and prove a lemma. We adopt the following labeling for vertices of $H\circ tK_{1}$ where $r$ is the order of $H$: Label the vertices of $H$ as $u_{0},\ldots,u_{r-1}$ and let the set of leaves adjacent to $u_{i}$ be denoted by $L_{i}=\{v_{i,j}:j=1,\ldots,t\}$ for $i=0,\ldots,r-1$. Furthermore, let $W_{i}=\{u_{i}\}\cup L_{i}$ for $i=0,\ldots,r-1$.
Lemma 5.10.
Let $H$ be a connected graph of order $r\geq 3$, let $t\geq 2$, let $G=H\circ tK_{1}$, and let $S$ be a ZIr-set of $G$. Then $G$ has a ZIr-set $S^{\prime}$ such that $|S|=|S^{\prime}|$ and every vertex of $S^{\prime}$ is a leaf.
Proof.
Let $S$ be a ZIr-set and let $S_{i}=S\cap W_{i}$. Note that if $F$ is a fort and $u_{i}\in F$, then $W_{i}\subseteq F$. So $u_{i}\in S$ implies $v_{i,j}\not\in S$ for $j=1,\dots,t$ because none of these vertices could have a private fort relative to $S$. If $u_{i}\in S$, we wish to replace $u_{i}$ by an element of $L_{i}$ in $S$ to obtain a ZIr-set of the same cardinality with one fewer vertex of $H$. Repeating this process as needed will establish the result. To replace $u_{i}$ by $v_{i,1}$, we need to verify that we can adjust other private forts if needed so that $v_{i,1}$ is not in the private fort of any element of $S$ except $u_{i}$. Let $x\in S_{k}$ with $k\neq i$. Choose a minimal private fort $F_{x}$. We show that minimality implies $|F_{x}\cap L_{i}|\leq 1$: Note that $u_{i}\not\in F_{x}$ since $u_{i}\in S$. The only reason for any $v_{i,j}$ to be in $F_{x}$ is that $u_{i}$ has exactly one neighbor in $F_{x}$ that is not in $L_{i}$. Thus including one leaf in $L_{i}$ is sufficient. If $F_{x}\cap L_{i}\neq\emptyset$, without loss of generality we modify $F_{x}$ so that $F_{x}\cap L_{i}=\{v_{i,t}\}$. Define $S^{\prime}=S\setminus\{u_{i}\}\cup\{v_{i,1}\}$. Then $S^{\prime}$ is a ZIr-set with private forts $\{v_{i,1},v_{i,t}\}$ for $v_{i,1}$ and $F_{x}$ for $x\not\in W_{i}$. ∎
Note that maximality need not be preserved in moving to an all-leaf ZIr-set of the same cardinality.
Theorem 5.11.
Let $r\geq 3$. Then
$\operatorname{Z}(C_{r}\circ 2K_{1})=\overline{\operatorname{Z}}(C_{r}\circ 2K_{1})=\operatorname{ZIR}(C_{r}\circ 2K_{1})=r=\frac{1}{3}|V(K_{r}\circ 2K_{1})|.$
Proof.
In $G=C_{r}\circ 2K_{1}$, we modify the labeling convention to label the two leaf neighbors of $u_{i}$ by $x_{i}$ and $y_{i}$. Index arithmetic is done in $\mathbb{Z}_{r}$, i.e., modulo $r$. Observe that $x_{i}$ and $y_{i}$ are twins for $i=1,\dots,r$, so at least one of them must be a zero forcing set. Thus $r\leq\operatorname{Z}(G)\leq\overline{\operatorname{Z}}(G)\leq\operatorname{ZIR}(G)$.
Now we show that $\operatorname{ZIR}(G)\leq r$. Let $W_{i}=\{u_{i},x_{i},y_{i}\}$. Let $S$ be a ZIr-set and let $S_{i}=S\cap W_{i}$. By Lemma 5.10, we may replace any occurrence of $u_{i}$ in $S$ by $x_{i}$ and assume $S$ consists entirely of leaves. Thus the possible values of $s=|S_{i}|$ are 00, $1$, or $2$. For $s=0,1,2$ define $T_{s}=\{i:|S_{i}|=s\}$ and observe that $r=|T_{2}|+|T_{1}|+|T_{0}|$ and $|S|=2|T_{2}|+|T_{1}|$. We show that $|T_{2}|\leq|T_{0}|$, which implies that $|S|\leq r$. For each $z\in S$ we choose a minimal private fort $F_{z}$. If $z\in S_{i}$, $u_{j}\in F_{z}$, and $j\neq i$, then $j\in T_{0}$ because $u_{j}\in F_{z}$ implies $W_{j}\subseteq F_{z}$.
If $T_{2}=\emptyset$, then there is nothing to prove, so without loss of generality (by renumbering if necessary) assume $0\in T_{2}$, so $S_{0}=\{x_{0},y_{0}\}$. Note that $y_{0}\not\in F_{x_{0}}$, which implies $u_{0}\notin F_{x_{0}}$. Thus $u_{0}$ needs at least two neighbors in $F_{x_{0}})$, so one of $u_{r-1}$ and $u_{1}$ must be in $F_{x_{0}}$. Without loss of generality, assume that $u_{1}\in F_{x_{0}}$, which implies $1\in T_{0}$. We proceed in increasing order around the cycle. If there are two consecutive cycle vertices $u_{j},u_{j+1}\not\in F_{x_{0}}$, then we stop as soon as this is encountered and set $j_{0}=j$. If we do not stop because there is no $j$ such that $u_{j},u_{j+1}\not\in F_{x_{0}}$, then at least half of the cycle vertices $u_{j}$ are in $F_{x_{0}}$. Thus $|T_{0}|\geq\frac{r}{2}\geq|T_{2}|$.
For the remainder of the proof we assume $F_{z}$ omits two consecutive cycle vertices for each $z\in S$. The assumption that $u_{1}\in F_{x_{0}}$ and we have stopped with $u_{j_{0}},u_{j_{0}+1}\not\in F_{x_{0}}$ and $u_{j_{0}-1}\in F_{x_{0}}$ implies at least one of $x_{j_{0}}$ or $y_{j_{0}}$ is in $F_{x_{0}}$. Define $I_{0}=\{0,1,\dots,j_{0}-1\}$. Observe that $|I_{0}\cap T_{0}|\geq|I_{0}\cap T_{2}|$ and $j_{0}\not\in T_{2}$.
We start with the interval $I=I_{0}$ and expand as needed through the process described next, such that the interval $I=[0,b]$ retains the following properties (which are true for $I_{0}$):
  1. 1.
    $|I\cap T_{0}|\geq|I\cap T_{2}|$.
  2. 2.
    If $b\in T_{2}$, then $b-1\in I\cap T_{0}$. If $b\in T_{0}$, then $b+1\not\in T_{2}$.
If $T_{2}\subset I$, then the proof is complete by condition (1). If not, let $p$ be the first element of $T_{2}$ that is not in $I$ starting at $b$ and proceeding in increasing order. We examine a private fort $F_{x_{p}}$. As in the construction of $I_{0}$, one of $u_{p-1}$ or $u_{p+1}$ is in $F_{x_{p}}$. If $u_{p-1}\in F_{x_{p}}$, then update the interval to $I=[0,\dots,p]$ and observe that the conditions (1) and (2) remain true because $p-1\in T_{0}$ and $j\not\in T_{2}$ for $j=b+1,\dots,p-2$ by the way $p$ was chosen. So suppose $u_{p-1}\not\in F_{x_{p}}$, which implies $u_{p+1}\in F_{x_{p}}$. We proceed in increasing order around the cycle. If for all $j=p+2,\dots,r-2$, $u_{j}\not\in F_{x_{p}}$ implies $u_{j+1}\in F_{x_{p}}$, then $I=\{0,\dots,r-1\}$ satisfies (1) and the proof is complete. So let $j$ be the first index (in increasing order) such that $u_{j},u_{j+1}\not\in F_{x_{p}}$. Then, update the interval to $I=[0,\dots,j-1]$ and observe that the conditions (1)–(2) remain true. This completes the proof. ∎
6 Effect of graph operations on upper ZIr number
In this section we examine the effect of graph operations in ZIr-sets and ZIr numbers, primarily the upper ZIr number. We first consider the effect of removing a cut-vertex.
Proposition 6.1.
Let $G$ be a connected graph such that $c$ is a cut-vertex of $G$ and $G-c$ has $\ell\geq 3$ components, denoted by $H_{1},\dots,H_{\ell}$ and ordered so that $\operatorname{ZIR}(H_{1})\geq\dots\geq\operatorname{ZIR}(H_{\ell})$. For $i=1,\dots,\ell$, let $S_{i}$ be a ZIr-set for $H_{i}$. Then $\bigcup_{i\neq k}S_{i}$ is a ZIr-set for $k=1,\dots,\ell$ and $\operatorname{ZIR}(G)\geq\operatorname{ZIR}(H_{1})+\dots+\operatorname{ZIR}(H_{\ell-1})$. This bound is sharp.
Proof.
Fix $k\in\{1,\dots,\ell\}$ and let $S=\bigcup_{i\neq k}S_{i}$. For each $x\in S$, there is an $i\neq k$ such that $x\in S_{i}$. Then $x$ has a private fort $F$ in $H_{i}$. If $|F\cap N(c)|=0$ or $|F\cap N(c)|\geq 2$, then $F$ is a private fort of $x$ in G relative to $S$. If $|F\cap N(c)|=1$, then $F^{\prime}=F\cup V(H_{k})$ is a private fort of $x$ in G relative to $S$. Since every $x\in S$ has a private fort of $x$ in G relative to $S$, $S$ is a ZIr-set, and by choosing $k=\ell$ we have $\operatorname{ZIR}(G)\geq\operatorname{ZIR}(H_{1})+\dots+\operatorname{ZIR}(H_{\ell-1})$.
To see that the bound is sharp, consider the graph $G=C_{r}\circ 2K_{1}$. Each cycle vertex is a cut-vertex with components $H_{1}\cong P_{r-1}\circ 2K_{1}$, $H_{2}\cong K_{1}$, and $H_{3}\cong K_{1}$. By Theorem 5.11, $\operatorname{ZIR}(C_{r}\circ 2K_{1})=r$. Furthermore, $\operatorname{Z}(P_{r-1}\circ 2K_{1})=r-1$ since $P_{r-1}\circ 2K_{1}$ is a tree. Thus $\operatorname{ZIR}(C_{r}\circ 2K_{1})=r\geq\operatorname{ZIR}(H_{1})+\operatorname{ZIR}(H_{2})=\operatorname{ZIR}(P_{r-1}\circ 2K_{1})+\operatorname{ZIR}(K_{1})\geq\operatorname{Z}(P_{r-1}\circ 2K_{1})+\operatorname{ZIR}(K_{1})=(r-1)+1=r$. ∎
We now consider the join of two graphs $G\vee H$. Proposition 6.2 handles the more general case and Proposition 6.3 is needed for the special case of $H=K_{1}$.
Proposition 6.2.
Let $G$ and $H$ be graphs on $n_{G}\geq 2$ and $n_{H}\geq 2$ vertices, respectively. Then $n_{G}+n_{H}-4\leq\operatorname{ZIR}(G\vee H)\leq n_{G}+n_{H}-1$. Both bounds are sharp. In particular, $\operatorname{ZIR}(P_{n_{G}}\vee P_{n_{H}})=n_{G}+n_{H}-4$.
Proof.
Let $D\subseteq V(G\vee H)$ such that $|D\cap V(G)|=2$ and $|D\cap V(H)|=2$. Then $D$ is a 2-dominating set. By Proposition 4.1, $\operatorname{ZIR}(G\vee H)\geq n_{G}+n_{H}-4$. Since $G\vee{H}$ has an edge, Remark 2.7 implies $\operatorname{ZIR}(G\vee H)\leq n_{G}+n_{H}-1$.
The upper bound is realized by $G\cong K_{n_{G}}$ and $H\cong K_{n_{H}}$. We now show that the lower bound is realized by $G\cong P_{n_{G}}$ and $H\cong P_{n_{H}}$ for $n_{G}\geq 7$ and $n_{H}\geq 7$. Let $v_{1},\ldots,v_{n_{G}}$ and $u_{1},\ldots,u_{n_{H}}$ denote the vertices of $G$ and $H$, respectively, where $v_{i}$ (respectively $u_{i}$) is adjacent to $v_{j}$ (respectively $u_{j}$) if and only if $i=j\pm 1$. Assume, by way of contradiction, that $\operatorname{ZIR}(G\vee H)\geq n_{G}+n_{H}-3$. Then $G\vee H$ has a ZIr-set $S$ of size $n_{G}+n_{H}-3$.
First, suppose $V(G)\subseteq S$. Observe that every private fort of every vertex in $S\cap V(H)$ relative to $S$ is a fort of $H$, and therefore contains $u_{1}$ and $u_{n_{H}}$. Thus, $u_{1},u_{n_{H}}\notin S$ and so there exists some $j\in\{2,n_{G}-1\}$ such that $u_{j}\in S$. This is a contradiction since every fort in $H$ containing $u_{j}$ has size at least 5, which implies $|S|\leq n_{G}+n_{H}-5+1$. Thus $V(G)$ is not a subset of $S$. Similarly, $V(H)$ is not a subset of $S$.
Without loss of generality, assume $|V(G)\setminus S|=1$ and $|V(H)\setminus S|=2$. Since $n_{H}\geq 7$, there exists some $u_{j}\in V(H)$ such that $N[u_{j}]\cap V(H)\subseteq S$, and there exists some $u_{k}\in S\cap V(H)$ such that $u_{k}$ is not adjacent to $u_{j}$ and $j\not=k$. Let $F_{u_{k}}$ be a private fort of $u_{k}$ relative to $S$. Since every fort of $H$ is a dominating set of $H$, $F_{u_{k}}$ is not a fort of $H$. Thus, $F_{u_{k}}\cap V(G)=\{v_{\ell}\}$ where $V(G)\setminus S=\{v_{\ell}\}$. This is a contradiction since $N[u_{j}]\cap F_{u_{k}}=\{v_{\ell}\}$. Thus, $\operatorname{ZIR}(G\vee H)=n_{G}+n_{H}-4$. ∎
The lower bound in the previous result can fail when $n_{G}=1$ and $n_{H}=2$, as illustrated by the friendship graph $\operatorname{Fr}(k)=K_{1}\vee kK_{2}$ since $|V(\operatorname{Fr}(k))|=2k+1$ and $\operatorname{ZIR}(\operatorname{Fr}(k))=k+1$ by Proposition 3.5.
Proposition 6.3.
Let $G$ be a graph of order $n_{G}$ with no isolated vertices. Then
$n_{G}-\gamma(G)\leq\operatorname{ZIR}(G\vee K_{1})\leq n_{G}-\gamma(G)+1.$ (6.1)
If $\operatorname{ZIR}(G)=n_{G}-\gamma(G)$, then $\operatorname{ZIR}(G\vee K_{1})=n_{G}-\gamma(G)+1=\operatorname{ZIR}(G)+1$. Both bounds in (6.1) are sharp.
Proof.
Let $u$ denote the vertex of $K_{1}$, let $S$ be a maximal ZIr-set of $G$, and let $D$ be a dominating set of $G$. Then $D\cup\{u\}$ is a 2-dominating set of $G\vee K_{1}$, so $\gamma_{2}(G\vee K_{1})\leq\gamma(G)+1$ and $\operatorname{ZIR}(G\vee K_{1})\geq(n_{G}+1)-(\gamma(G)+1)=n_{G}-\gamma(G)$ by Proposition 4.1.
Let $\hat{S}$ be a $\operatorname{ZIR}$-set of $G\vee K_{1}$. If $u$ is in a private fort $F$ of some vertex in $\hat{S}$, then $F$ contains at least $\gamma(G)$ vertices of $G$ since $u$ is adjacent to every vertex in $G$. Therefore $|\hat{S}|\leq(|V(G)|+1)-|F|+1=n_{G}+1-(\gamma(G)+1)+1=n_{G}-\gamma(G)+1$. If $u$ is not in any private fort of a vertex in $\hat{S}$, then $u\not\in\hat{S}$ and every private fort relative to $\hat{S}$ in $G\vee K_{1}$ is also a private fort relative to $\hat{S}$ in $G$. Therefore $\hat{S}$ is a ZIr-set of $G$, and $\operatorname{ZIR}(G\vee K_{1})=|\hat{S}|\leq\operatorname{ZIR}(G)\leq n_{G}-\gamma(G)$ by Proposition 4.1. Thus in either case $\operatorname{ZIR}(G\vee K_{1})\leq n_{G}-\gamma(G)+1$.
By Lemma 2.9, $V(G)\setminus S$ is a dominating set of $G$ (recall $S$ is a maximal ZIr-set of $G$). Thus, $\{u\}\cup(V(G)\setminus S)$ is a private fort of $u$ in $G\vee K_{1}$ relative to $S\cup\{u\}$. So $\operatorname{ZIR}(G\vee K_{1})\geq\operatorname{ZIR}(G)+1$. Thus $\operatorname{ZIR}(G)=n_{G}-\gamma(G)$ implies $\operatorname{ZIR}(G\vee K_{1})\geq n_{G}-\gamma(G)+1$.
The complete graph $G=K_{n}$ shows the upper bound is sharp since $n_{G}=n$, $\gamma(K_{n})=1$, $K_{n}\vee K_{1}\cong K_{n+1}$, and $\operatorname{ZIR}(K_{n+1})=n$. By Corollary 5.3, the graph $C_{r}\vee K_{2}\cong(C_{r}\vee K_{1})\vee K_{1}$ has $\operatorname{ZIR}(C_{r}\vee K_{2})=r=(r+1)-1=|V(C_{r}\vee K_{1})|-\gamma(C_{r}\vee K_{1})$. ∎
The wheel graph of order $r+1$ is $W_{r+1}=C_{r}\vee K_{1}$.
Proposition 6.4.
Let $r\geq 5$. Then $\operatorname{ZIR}(W_{r+1})=r-\gamma(C_{r})=r-\left\lceil\frac{r}{3}\right\rceil$ and $\operatorname{zir}(W_{r+1})=\operatorname{Z}(W_{r+1})=\overline{\operatorname{Z}}(W_{r+1})=3$.
Proof.
To establish $\operatorname{ZIR}(W_{r+1})=r-\gamma(C_{r})$, it suffices to show that $\operatorname{ZIR}(W_{r+1})\leq r-\gamma(C_{r})$ by Proposition 6.3. Let $u$ denote the vertex of $K_{1}$, let $\hat{S}$ be a maximal ZIr-set of $W_{r+1}$, and for $x\in\hat{S}$, let $F_{x}$ denote private fort of $x$ with respect to $\hat{S}$ in $W_{r+1}$.
Suppose first that $u\not\in F_{y}$ for some $y\in\hat{S}\cap V(C_{r})$. Then $F_{y}$ is a private fort of $y$ in $C_{r}$ relative to $\hat{S}\setminus\{u\}$. Note that every fort of $C_{r}$ contains at least $\left\lceil\frac{r}{2}\right\rceil$ vertices. If $u\not\in\hat{S}$, then $|\hat{S}|\leq r-\left\lceil\frac{r}{2}\right\rceil+1=\left\lfloor\frac{r}{2}\right\rfloor+1$. If $u\in\hat{S}$, then $u\not\in F_{y}$ for every $y\in\hat{S}\cap C_{r}$. Thus $\hat{S}\setminus\{u\}$ is a ZIr-set of $C_{r}$ and $|\hat{S}|\leq\operatorname{ZIR}(C_{r})+1=\left\lfloor\frac{r}{2}\right\rfloor+1$. Thus in both cases $u\not\in\hat{S}$ and $u\in\hat{S}$, $|\hat{S}|\leq\left\lfloor\frac{r}{2}\right\rfloor+1\leq r-\left\lceil\frac{r}{3}\right\rceil$ for $r\geq 5$, where the last inequality is verified algebraically.
Let $\hat{S}\cap V(C_{r})=\{y_{1},\dots,y_{\ell}\}$ and suppose that $u\in F_{y_{i}}$ for $i=1,\dots,\ell$ (which implies $u\not\in\hat{S}$). It suffices to show that $\ell\leq r-\gamma(C_{r})$. Since $u\in F_{y_{i}}$, $F_{y_{i}}$ must contain a dominating set for $C_{r}$ for every $y_{i}$. Recall that $|\hat{S}|\leq r+1-|\cup_{i=1}^{k}F_{y_{i}}|+k=r-|\left(\cup_{i=1}^{k}F_{y_{i}}\right)\cap V(C_{r})|+k$ for $k=1,\dots,\ell$ (Observation 2.13). In order to have $|\hat{S}|=r-|\left(\cup_{i=1}^{k}F_{y_{i}}\right)\cap V(C_{r})|+k$ there must be a set $\hat{D}$ such that $\hat{D}\cup\{y_{i}\}$ is a minimum dominating set for $i=1,\dots,\ell$. Since $r\geq 5$, there does not exist such a set for $\ell=r-\gamma(C_{r})$.
It is well-known that $\operatorname{Z}(W_{r+1})=3$ and easy to see that $\overline{\operatorname{Z}}(W_{r+1})=3$ since any zero forcing set must contain two consecutive cycle vertices and $u$, or three consecutive cycle vertices, and any set of this type is minimal. By Corollary 2.11, $3=\delta(W_{r+1})=3\leq\operatorname{zir}(W_{r+1})$. ∎
We now consider the corona product of two graphs. The following notational conventions will be used throughout the rest of this section. Let $G$ and $H$ be graphs on $n_{G}$ and $n_{H}$ vertices, respectively. Let $v_{1},\ldots,v_{n_{G}}$ denote the vertices of $G$. For $i=1,\ldots,n_{G}$ let $H_{i}$ denote the copy of $H$ in $G\circ H$ whose vertices are adjacent to $v_{i}$, and let $H_{i}^{\prime}$ denote the subgraph of $G\circ H$ induced by $V(H_{i})\cup\{v_{i}\}$, which is isomorphic to $H\vee K_{1}$. The next result states basic facts about forts and ZIr-sets of coronas $G\circ H$ when $H$ has no isolated vertices.
Proposition 6.5.
Let $G$ be a graph, let $H$ be a graph with no isolated vertices, and let $F$ be a fort of $G\circ H$. Then $F^{\prime}_{i}=F\cap V(H^{\prime}_{i})$ is the empty set or a fort of $H^{\prime}_{i}$. If $S$ is a ZIr-set of $G\circ H$, then $S^{\prime}_{i}=S\cap V(H_{i}^{\prime})$ is a ZIr-set of $H_{i}^{\prime}$ and $S=\cup_{i=1}^{n_{G}}S^{\prime}_{i}$. Thus $\operatorname{ZIR}(G\circ H)\leq n_{G}\operatorname{ZIR}(H\vee K_{1})$.
Proof.
Since $H$ has no isolated vertices, every fort of $H$ has at least two elements. Thus a fort of $H_{i}$ is also a fort of $H^{\prime}_{i}$. Assume $F^{\prime}_{i}\neq\emptyset$. If $v_{i}\in F^{\prime}_{i}$, then $F^{\prime}_{i}$ is a fort of $H^{\prime}_{i}$. If $v_{i}\not\in F^{\prime}_{i}$, then $F^{\prime}_{i}$ is a fort of $H_{i}$, so is also a fort of $H^{\prime}_{i}$. The remaining statements follow from this property of forts. ∎
Before proving our first result on the corona product of two graphs we introduce some new terminology. Let $S_{0}$ be a maximal ZIr-set of a graph $G$. An abandoned fort relative $S_{0}$ is a fort in $G$ that contains no elements of $S_{0}$ and $S_{0}$ abandons a fort if there exists an abandoned fort relative to $S_{0}$. We say that $G$ abandons a fort if there exists an upper ZIR set $S$ that abandons a fort.
Example 6.6.
Many abandoned forts naturally occur as a result of 2-dominating sets. This is not surprising since a fort is a 2-dominating set of its closed neighborhood. Examples we have already considered include $P_{n},n\geq 3$ and $C_{n},n\geq 4$. The two vertices not in $H$ are an abandoned fort relative to $S=V(H)$ for $H\vee 2K_{1}$, or $H\vee K_{2}$ when $H\not\cong K_{r}$ (cf. Corollary 5.3). Figure 6.1 provides an example of a connected graph that abandons a fort that is not a 2-dominating set: the upper ZIR set $\{v_{1},v_{4},v_{5}\}$ abandons the fort $\{v_{2},v_{3}\}$, which is not 2-dominating.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$
Figure 6.1: A graph with an upper ZIR-set that abandons a fort that is not 2-dominating.
A single vertex forms a fort if and only if that vertex is isolated. Thus, we have the following observation.
Observation 6.7.
If a maximal ZIr-set abandons a fort $F$, then $|F|\geq 2$.
Theorem 6.8.
Let $G$ and $H$ be graphs on $n_{G}$ and $n_{H}$ vertices, respectively. Assume $H$ has no isolated vertices and $H\vee K_{1}$ abandons a fort. Then $\operatorname{ZIR}(G\circ H)=n_{G}\operatorname{ZIR}(H\vee K_{1})$.
Proof.
By Remark 6.5, $\operatorname{ZIR}(G\circ H)\leq n_{G}\operatorname{ZIR}(H\vee K_{1})$. Equality is obtained by constructing an appropriately sized ZIr-set.
Let $S=S_{1}\cup\cdots\cup S_{n_{G}}$, where each $S_{i}$ is an upper ZIR set of $H_{i}^{\prime}$ that abandons a fort $F_{i}$. Without loss of generality assume that either $v_{i}\in F_{i}$ for every $i=1,\ldots,n_{G}$, or $v_{i}\notin F_{i}$ for every $i=1,\ldots,n_{G}$. Let $u\in S_{k}$ for some $k\in\{1,\ldots,n_{G}\}$ and let $F$ be a private fort of $u$ relative to $S_{k}$ in $H_{k}^{\prime}$. It suffices to show $u$ has a private fort relative to $S$ in $G\circ H$.
First, suppose $v_{k}\in F$. If $v_{i}\in F_{i}$ for every $i=1,\ldots,s$, then the union of $F$ with every $F_{i}\not=F_{k}$ is a private fort of $u$ relative to $S$ in $G\circ H$. If $v_{i}\notin F_{i}$ for every $i=1,\ldots,s$, then the union of $F$ with each $F_{i}$ such that $v_{i}\in N(v_{k})$ is a private fort of $u$ relative to $S$ in $G\circ H$.
Now suppose $v_{k}\notin F$. Then every vertex in $V(G\circ H)\setminus V(H_{k}^{\prime})$ is adjacent to no vertices in $F$. Thus, $F$ is a private fort of $u$ relative to $S$ in $G\circ H$. Thus $S$ is a ZIr-set of $G\circ H$ and so $\operatorname{ZIR}(G\circ H)\geq|S|=n_{G}\operatorname{ZIR}(H\vee K_{1})$. ∎
Remark 6.9.
Let $G$ be a graph and let $S$ be a maximal ZIr-set set of $G$. Then $S$ abandons a fort if and only if $S$ is not a zero forcing set, because $S$ is a zero forcing set if and only if it intersects every fort by Theorem 1.1. Furthermore, if $G$ does not abandon a fort, then every upper ZIR set is a zero forcing set, so $\operatorname{ZIR}(G)=\overline{\operatorname{Z}}(G)$. Thus Theorem 6.8 applies to any graph $H$ with no isolated vertices such that $\overline{\operatorname{Z}}(H\vee K_{1})<\operatorname{ZIR}(H\vee K_{1})$.
The next two examples provide families of graphs $H$ to which the previous theorem applies.
Example 6.10.
If $\operatorname{ZIR}(H\vee K_{1})=n_{H}-\gamma(H)$ and $H\not\cong K_{r}$, then $H\vee K_{1}$ abandons a fort: Let $D_{H}$ be a dominating set of $H$ and let $u$ be the vertex of $K_{1}$. Then the 2-dominating set $F=D_{H}\cup\{u\}$ is an abandoned fort of the upper ZIR set $S=V(H)\setminus D_{H}$ for $H\vee K_{1}$. In particular, $W_{r+1}=C_{r}\vee K_{1}$ and $\operatorname{ZIR}(W_{r+1})=|V(C_{r})|-\gamma(C_{r})$ by Proposition 6.4. Thus $\operatorname{ZIR}(G\circ C_{r})=n_{G}(r-\left\lceil\frac{r}{3}\right\rceil)$ for $r\geq 4$.
Example 6.11.
Let $\hat{H}$ be a graph that is not complete. Define $H=\hat{H}\vee K_{1}$. Since $\hat{H}$ is not complete, neither is $H\vee K_{1}$, so $\operatorname{ZIR}(H\vee K_{1})\leq|V(H\vee K_{1})|-2=|V(\hat{H})|$. As noted in Example 6.6, $F=V(H\vee K_{1})\setminus V(\hat{H})$ is an abandoned fort relative to the upper ZIR set $S=V(\hat{H})$.
Next, we present an example of a graph $H$ such that $H$ does not abandon a fort but $H\vee K_{1}$ does abandon a fort.
Example 6.12.
Recall that $\operatorname{Fr}(k)=K_{1}\vee kK_{2}$ denotes the $k$th friendship graph. It was shown in Proposition 3.5 that any upper ZIR set $S$ of $\operatorname{Fr}(k)$ must contain $v_{0}$ (the vertex of degree $2k$. Any fort $F_{v_{0}}$ of $\operatorname{Fr}(k)$ that contains $v_{0}$ also contains at least one vertex from each $K_{2}$. However, $S$ must contain a vertex from every $K_{2}$ or it would not be maximal. Since every fort of $\operatorname{Fr}(k)$ contains $v_{0}$ or $S$ contains either 0 or 2 for each $K_{2}$, $\operatorname{Fr}(k)$ does not abandon a fort. Since $\operatorname{Fr}(k)\vee K_{1}\cong 2K_{2}\vee K_{2}$, $\operatorname{Fr}(k)\vee K_{1}$ abandons a fort.
We now analyze $G\circ H$ when $H\vee K_{1}$ does not abandon a fort. Recall that when $H$ has no isolated vertices, every fort of $H$ contains at least two vertices, which implies the next observation.
Observation 6.13.
If $H$ contains no isolated vertices, then every fort of $H_{i}$ is a fort of $G\circ H$.
Observation 6.14.
If $n_{H}\geq 2$, then $\operatorname{ZIR}(G\circ H)\geq n_{G}$ because $V(H_{i})$ is a fort of $G\circ H$ for $i=1,\dots,n_{G}$. This bound is sharp since $\operatorname{ZIR}(C_{r}\circ 2K_{1})=r$ by Theorem 5.11.
Our next result is a lower bound that applies when assuming only that $H$ has no isolated vertices (although Theorem 6.8 determines $\operatorname{ZIR}(G\circ H)$ when $H\vee K_{1}$ abandons a fort).
Proposition 6.15.
Let $G$ and $H$ be graphs on $n_{G}$ and $n_{H}$ vertices, respectively. If $H$ contains no isolated vertices, then $\operatorname{ZIR}(G\circ H)\geq\operatorname{ZIR}(G)\operatorname{ZIR}(H\vee K_{1})+(n_{G}-\operatorname{ZIR}(G))\operatorname{ZIR}(H)$. The bound is sharp.
Proof.
Let $R$ be a ZIr-set of $G$ such that $|R|=\operatorname{ZIR}(G)$. Without loss of generality $R=\{v_{1},\ldots,v_{k}\}$, where $k=\operatorname{ZIR}(G)$. For $i=1,\ldots,k$ let $F_{i}$ denote a private fort of $v_{i}$ relative to $S$ in $G$, and let $S_{i}$ be an upper ZIR set of $H_{i}^{\prime}$. For $i=k+1,\ldots,{n_{G}}$ let $S_{i}$ be an upper ZIR set of $H_{i}$, and let $T_{i}=V(H_{i})\setminus S_{i}$. Let $S=S_{1}\cup\cdots\cup S_{n_{G}}$ and let $u\in S$. Then $u\in S_{j}$ for some $j\in\{1,\dots,{n_{G}}\}$.
First, assume $j\in\{k+1,\dots,{n_{G}}\}$. Then $u$ has a private fort $F$ relative to $S_{j}$ in $H_{j}$. By Observation 6.13, $F$ is a private fort of $u$ relative to $S$ in $G\circ H$.
Now assume $j\in\{1,\ldots,k\}$. Then $u$ has a private fort $F$ relative to $S_{j}$ in $H_{j}^{\prime}$. If $v_{j}\notin F$, then $F$ is a private fort of $u$ relative to $S$ in $G\circ H$. So, suppose $v_{j}\in F$ and let $T$ be the union of each $T_{i}$ such that $v_{i}\in F_{j}$. Then $F\cup F_{j}\cup T$ is a private fort of $u$ relative to $S$ in $G\circ H$. Thus, $S$ is a ZIr-set and $\operatorname{ZIR}(G\circ H)\geq|S|=\operatorname{ZIR}(G)\operatorname{ZIR}(H\vee K_{1})+(n_{G}-\operatorname{ZIR}(G))\operatorname{ZIR}(H)$.
When $G=K_{1}$, $G\circ H=H\vee K_{1}$ for any $H$. Thus $\operatorname{ZIR}(G)\operatorname{ZIR}(H\vee K_{1})+(n_{G}-\operatorname{ZIR}(G))\operatorname{ZIR}(H)=(1)\operatorname{ZIR}(H\vee K_{1})+(1-1)\operatorname{ZIR}(H)=\operatorname{ZIR}(H\vee K_{1})=\operatorname{ZIR}(G\circ H)$. ∎
Although Proposition 6.15 excluded isolated vertices for $H$, it is easy to compute $\operatorname{ZIR}(G\circ K_{1})$, as in the next example.
Example 6.16.
Let $G$ be a graph of order $r$. We can see that the set $S=V(G)$ is a ZIr-set of cardinality $r$: Let $u_{i}\in V(H_{i})$ be the unique neighbor of $v_{i}$ that is not in $V(G)$. Then $F_{i}=\{v_{i}\}\cup\{u_{j}:v_{j}\in N[v_{i}]\}$ is a private fort of $v_{i}$. By Proposition 4.1 $\operatorname{ZIR}(G\circ K_{1})\leq 2r-\gamma(G\circ K_{1})=2r-r=r$, so $\operatorname{ZIR}(G\circ K_{1})=r$. It is known (and easy to see, e.g., by considering consider $C_{4k}$ vs. $K_{r}$) that $\operatorname{Z}(G\circ K_{1})$ depends on $G$.
A set of vertices $A$ is independent if no two vertices in $A$ are adjacent. The independence number of a graph $G$, denoted by $\alpha(G)$, is the size of a maximum independent set for $G$. The next result may be useful in the case where $H\vee K_{1}$ does not abandon a fort and $H$ has one or more isolated vertices.
Proposition 6.17.
Let $G$ be a graph on $n_{G}$ vertices and let $H$ be a graph. Then $\operatorname{ZIR}(G\circ H)\geq\alpha(G)\operatorname{ZIR}(H\vee K_{1})+(n_{G}-\alpha(G))(\operatorname{ZIR}(H)-1)$ and this bound is sharp.
Proof.
Let $A$ be an independent set of $G$ such that $|A|=\alpha(G)$. For each $v_{i}\in A$, let $S^{\prime}_{i}$ be an upper ZIR set of $H_{i}^{\prime}$. For each $v_{i}\in V(G)\setminus A$, let $R_{i}$ be an upper ZIR set of $H_{i}$. Let $u_{i}\in R_{i}$ with private fort $F_{i}$ relative to $R_{i}$ in $H_{i}$. Then $S_{i}=R_{i}\setminus\{u_{i}\}$ abandons the fort $F_{i}$ in $H_{i}$.
Let $S=(\bigcup_{i:v_{i}\in A}S_{i}^{\prime})\cup(\bigcup_{i:v_{i}\notin A}S_{i})$. We show that $S$ is a ZIr-set of $G\circ H$. Let $u\in S$. Suppose first that $u\in{S_{j}}$ for some $j=1,\ldots,n_{G}$. Then $u$ has a private fort $F$ relative to $S_{j}$ in $H_{j}$ and hence, $F\cup F_{j}\subseteq V(H_{j})$ is a private fort of $u$ relative to $S_{j}$ in $H_{j}^{\prime}$ because $|F\cup F_{j}|\geq 2$. Now suppose $u\in S^{\prime}_{j}$ for some $j=1,\ldots,n_{G}$. If $u$ has a private fort $F\subseteq V(H_{j})$ relative to $S\cap V(H_{j}^{\prime})$ in $H_{j}^{\prime}$, then $F$ is a private fort of $u$ relative to $S$ in $G\circ H$. So, suppose this is not the case. Then $u$ has a private fort $F$ relative to $S\cap V(H_{j}^{\prime})$ that contains $v_{j}$. Then the union of $F$ with every $F_{i}$ such that $v_{i}\in N(v_{j})$ is a private fort of $u$ relative to $S$ in $G\circ H$. Thus, $S$ is a ZIr-set and $\operatorname{ZIR}(G\circ H)\geq\alpha(G)\operatorname{ZIR}(H\vee K_{1})+(n_{G}-\alpha(G))(\operatorname{ZIR}(H)-1)$.
For sharpness, consider any graph $G$ of order $n_{G}$ and $H=sK_{1}$ with $s>n_{G}$. Then $\operatorname{ZIR}(H)=s$ and $\operatorname{ZIR}(H\vee K_{1})=\operatorname{ZIR}(K_{1,s})=s-1$. It follows that $\operatorname{ZIR}(G\circ H)\geq\alpha(G)(s-1)+(n_{G}-\alpha(G))(s-1)=n_{G}\operatorname{ZIR}(H\vee K_{1})=n_{G}(s-1)$. Let $S$ be an upper ZIR set of $G\circ H$. Note that if $v_{i}\in S$, then $S\cap V(H_{i}^{\prime})=\{v_{i}\}$ and therefore $|S\cap V(H_{i}^{\prime})|\leq s$. If there exist an $H_{i}^{\prime}$ such that $|S\cap V(H_{i}^{\prime})|=s$, then $V(H_{i})\subseteq S$ and $v_{i}$ is in the private fort $F_{v}$ of every $v\in S\cap H_{i}$. To ensure that $v_{i}$ is adjacent to two vertices of $F_{v}$, there exist a neighbour of $v_{i}$ in $G$ that is in $F_{v}$, say $v_{k}$. Since $v_{k}\in F_{v}$ it follows that $V(H_{k})\subseteq F_{v}$. Therefore $|S|\leq(n_{G}-1)s=sn_{G}-s$ and since $s>n_{G}$ it contradicts the lower bound on $\operatorname{ZIR}(G\circ H)$. It follows that $|S\cap V(H_{i})|\leq s-1$ for each $i$ and therefore $\operatorname{ZIR}(G\circ H)=n_{G}(s-1)$. ∎
7 Nonrelationships
In this section we present examples that show that some parameters that are lower bounds for zero forcing number are noncomparable to the lower ZIr number. We also show that the independence number is noncomparable to $\operatorname{zir}$ and $\operatorname{ZIR}$; this is not surprising since it is well-known that it is not comparable to $\operatorname{Z}$, but worth noting since $\alpha$ appears in a lower bound for $\operatorname{ZIR}$.
One of the origins of zero forcing of a graph $G$ was the study of maximum nullity among the set of symmetric matrices that have graph $G$. Let $G$ be a graph with $V(G)=\{v_{1},\dots,v_{n}\}$. The graph of a real symmetric $n\times n$ matrix $A$, denoted by $\mathcal{G}(A)$, has vertex set $\{v_{1},\dots,v_{n}\}$ and has edge set $\{v_{i}v_{j}:i\neq j\mbox{ and }a_{ij}\neq 0\}$. The maximum nullity of the graph $G$, denoted by $\operatorname{M}(G)$, is defined to be
$\operatorname{M}(G)=\max\{\operatorname{null}A:\mathcal{G}(A)=G\mbox{ and }A^{T}=A\}$
where $A^{T}$ denotes the transpose of $A$. It is known that $\operatorname{M}(G)\leq\operatorname{Z}(G)\leq\operatorname{ZIR}(G)$ for all $G$, so it seems natural to ask whether there is a relationship between maximum nullity and lower ZIR number. The next examples show there is no relationship between $\operatorname{M}$ and $\operatorname{zir}$.
Example 7.1.
The star $K_{1,p}$ has $\operatorname{zir}(K_{1,p})=1<p-1=\operatorname{M}(K_{1,p})$ by Example 3.1 (maximum nullity of the star is well-known [15, Theorem 9.5]).
Example 7.2.
Let $G$ be the pentasun shown in Figure 7.1. Then it can be verified computationally that $\operatorname{zir}(G)=3$ [13] and it is known that $\operatorname{M}(G)=2$ [15, Example 9.118].
Figure 7.1: A graph $G$ having $M(G)<\operatorname{zir}(G)$
For any positive integer $k$, a $k$-tree is $K_{k+1}$ or a graph that is obtained from $K_{k+1}$ by repeatedly adding a new vertex and joining it to an existing $k$-clique. For a graph $G$, the minimum $k$ such that $G$ is a subgraph of some $k$-tree is called the tree-width of $G$, denoted by $\operatorname{tw}(G)$. It is known that $\operatorname{tw}(G)\leq\operatorname{Z}(G)$ for all $G$ [15], so it seems natural to ask whether there is a relationship between tree-width and lower ZIR number. The next examples show tree-width and lower ZIR number are not comparable.
Example 7.3.
As shown in Proposition 3.5, the friendship graph $\operatorname{Fr}(k)$ has $\operatorname{zir}(\operatorname{Fr}(k))=k+1$. To see that $\operatorname{tw}(\operatorname{Fr}(k))=2$, construct $\operatorname{Fr}(k)$ as a subgraph of a 2-tree $G$: First define a sequence of 2-trees $L_{i}$ as follows: $L_{1}=K_{3}$ with $V(L_{1})=\{0,1,2\}$. For $i=1,\dots,2k-2$, construct $L_{i+1}$ by adding vertex $i+2$ adjacent to exactly vertices 00 and $i+1$ of $L_{i}$. Observe that $V(L_{i})=\{0,1,\dots,i+1\}$ and $L_{i}$ is a $2$ tree. From $L_{2k-1}$, delete the edges $\{\{2i,2i+1\}:i=1,\dots,k-1\}$ to obtain $\operatorname{Fr}(k)$.
Example 7.4.
Let $G$ be the graph shown in Figure 7.2. Then it has been verified computationally that $\operatorname{zir}(G)=4$ and $\operatorname{tw}(G)=5$. [13].
Figure 7.2: A graph $G$ having $\operatorname{tw}(G)>\operatorname{zir}(G)$
Since $\gamma_{P}(G)\leq\operatorname{Z}(G)$, it is immediate that $\gamma_{P}(G)\leq\operatorname{ZIR}(G)$. Note that $C_{r}\circ 2K_{1}$ is an extremal example for power domination (least possible value of $\gamma_{P}(G)$ for graphs with $3r$ vertices) and may be an extremal example for $\operatorname{ZIR}(G)$ (may be least possible value of $\operatorname{ZIR}(G)$ for graphs with $3r$ vertices).
The next example shows that not every maximal ZIr-set is a power dominating set, but does not show that $\operatorname{zir}(G)<\gamma_{P}(G)$ is possible.
Example 7.5.
The graph $G$ in Figure 7.3 satisfies $\operatorname{zir}(G)=2$ and $\gamma_{P}(G)=2$. Observe that $S=\{v_{3},v_{4}\}$ is a maximal ZIr-set of $G$. However, $S$ is not a power dominating set.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$
Figure 7.3: A graph with a lower zir set that is not power dominating.
Question 7.6.
Is $\gamma_{P}(G)\leq\operatorname{zir}(G)$ for all graphs $G$?
Well-known examples that show $\alpha$ is noncomparable with $\operatorname{Z}$ also show that $\alpha$ is noncomparable with $\operatorname{zir}$ and $\operatorname{ZIR}$.
Example 7.7.
Consider $K_{n}$ with $n\geq 3$: $\alpha(K_{n})=1<n-1=\operatorname{zir}(K_{n})$.
Example 7.8.
Consider $C_{r}\circ 2K_{1}$ with $r\geq 3$: $\alpha(C_{r}\circ 2K_{1})=2r>r=\operatorname{ZIR}(C_{r}\circ 2K_{1})$.
The domination number has appeared in several bounds for $\operatorname{ZIR}$. It is known that $\gamma$ is noncomparable with $\operatorname{Z}$. Consider, for example, $K_{n},n\geq 3$ ($\gamma(K_{n})=1$ and $\operatorname{Z}(K_{n})=n-1$) and $C_{r}\circ K_{1}$ where $r=4\ell$ ($\gamma(C_{r}\circ K_{1})=r$ and $\operatorname{Z}(C_{r}\circ K_{1})=\frac{r}{2}$). What can be said $\gamma$ and $\operatorname{ZIR}$? From Corollary 4.3 a graph $G$ with $\delta(G)\geq 3$ has $\operatorname{ZIR}(G)\geq\frac{n}{2}$. Since a graph $G$ with no isolated vertices has $\gamma(G)\leq\frac{n}{2}$, it follows that $\gamma(G)\leq\operatorname{ZIR}(G)$ for graphs with minimum degree at least 3.
Question 7.9.
Is $\gamma(G)\leq\operatorname{ZIR}(G)$ for all graphs $G$?
8 Concluding remarks
In this section we summarize the values of lower ZIr number, zero forcing number, upper zero forcing number, and upper ZIr number for various graph families in Table 8.1 and present some open questions for future work. In most cases listed in Table 8.1, the zero forcing number was already known and most can be found (with references) in [15, Theorem 9.5].
result # graph $G$ order $\operatorname{zir}(G)$ $\operatorname{Z}(G)$ $\overline{\operatorname{Z}}(G)$ $\operatorname{ZIR}(G)$
2.6 $\overline{K_{n}}$ $n$ $n$ $n$ $n$ $n$
2.12 $K_{n}$ $n$ $n-1$ $n-1$ $n-1$ $n-1$
[…]
3.3 $P_{n}$,   $5\leq n$ $n$ $1$ $1$ $2$ $\left\lfloor\frac{n-1}{2}\right\rfloor$
[…]
3.5 $\operatorname{Fr}(k),\ k\geq 2$ $12k+$ 1 $k+1$ $k+1$ $k+1$ $k+1$
3.6 $H(r,s),\,r\geq 2,s\geq 5,s$ odd $r+s+1$ 2 $r$ $r+1$ $r+\frac{s-1}{2}$
$H(r,3),\,r\geq 2$ $r+s+1$ 2 $r$ $r$ $r+1$
4.4, 4.5 $N_{k}$ $4k$ $k+2$ $k+2$ $2k$
4.8 $H_{k}$ $5k$ $k+2$ $2k$
[…]
5.6 $rK_{2}\vee 2K_{1}$ $22r+$ 2 $r+2$ $2r$
6.2 $P_{n_{1}}\vee P_{n_{2}},\ n_{1},n_{2}\geq 7$ $n_{1}+n_{2}$ $n_{1}+n_{2}-4$
6.4 $W_{r+1}=C_{r}\vee K_{1},\,r\geq 5$ $r+1$ $3$ $3$ $3$ $r-\left\lceil\frac{r}{3}\right\rceil$
5.3 $H\vee 2K_{1}$ $|V(H)|+2$ $|V(H)|$
5.3 $H\vee K_{2},\,H\neq K_{|V(H)|}$ for $\operatorname{ZIR}$, $|V(H)|+2$ $\operatorname{Z}(H)+2$ $\overline{\operatorname{Z}}(H)+2$ $|V(H)|$
$\delta(H)\geq 1$ for $\operatorname{Z},\overline{\operatorname{Z}}$
6.10 $H\circ C_{r},\,r\geq 5$ $|V(H)|(r+1)$ $|V(H)|\left(r-\left\lceil\frac{r}{3}\right\rceil\right)$
6.11 $H\circ W_{r+1},\,r\geq 5$ $|V(H)|(r+2)$ $|V(H)|r$
6.16 $H\circ K_{1}$ $2|V(H)|$ $|V(H)|$
Table 8.1: Summary of values of $\operatorname{zir}(G),\operatorname{Z}(G),\overline{\operatorname{Z}}(G),$ $\operatorname{ZIR}(G)$ for various graph families $G$
Next, we list several questions that were implicit in earlier discussions.
Question 8.1.
What is the lowest value of $\operatorname{ZIR}(G)$ over all graphs of order $n$, as $n\to\infty$?
Question 8.2.
Theorem 5.11 provides a family of graphs $G$ satisfying $\operatorname{ZIR}(G)=\frac{1}{3}|V(G)|$, but these graphs have $\delta(G)=1$. What is an asymptotically tight lower bound on $\operatorname{ZIR}(G)$ for connected graphs $G$ of order $n$ with $\delta(G)=2$ as $n\to\infty$?
Question 8.3.
If $G$ is a connected cubic graph of order $n\geq 4$, then $\operatorname{ZIR}(G)\leq\frac{3}{4}n$ by Corollary 4.11, and equality is realized by $K_{4}$. Do there exist connected cubic graphs $G$ of arbitrarily large order $n$ such that $\operatorname{ZIR}(G)=\frac{3}{4}n$? If not, what is an asymptotically tight upper bound on $\operatorname{ZIR}(G)$ for connected cubic graphs $G$ of order $n$ as $n\to\infty$?
Section 6 only scratches the surface of how graph operations affect ZIr numbers. There are many related questions remaining. For example, what is the effect of graph operations on the lower zir number?
Question 8.4.
If $H$ abandons a fort, is it necessarily the case that $H\vee K_{1}$ abandons a fort?
Section 7 contains two questions about whether the lower zir number is at least the power domination number and whether the upper ZIR number is at least the domination number. We can also look for relationships and non-relationships with additional parameters.
Cockayne and Mynhardt [12] determined conditions under which integers $1\leq k_{1}\leq k_{2}\leq k_{3}\leq k_{4}$ can be realized as $\operatorname{ir}(G)=k_{1},\gamma(G)=k_{2},\Gamma(G)=k_{3}$, and $\operatorname{IR}(G)=k_{4}$. One condition is that $k_{1}=1$ implies that $k_{2}=1$; this condition would however not be true for the ZIr-numbers and zero forcing numbers since $\operatorname{zir}(K_{1,r})=1$, but $\operatorname{Z}(K_{1,r})=r-1$. We know from Proposition 3.6 that $k_{1}=2,k_{2}\geq k_{1}$, $k_{3}=k_{2}+1$, and $k_{4}>k_{3}$ can be realized by $H(r,s)$.
Question 8.5.
For which integer values $1\leq k_{1}\leq k_{2}\leq k_{3}\leq k_{4}$ does there exist a graph such that $\operatorname{zir}(G)=k_{1},Z(G)=k_{2},\bar{Z}(G)=k_{3}$ and $\operatorname{ZIR}(G)=k_{4}$?
Acknowledgements
The research of B. Curtis is partially supported by NSF grant 1839918. The research of A. Roux is supported in part by the National Research Foundation of South Africa (Grant Number: 121931).